### v1

# Spectrum radii of trees

Abstract
For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of positive numbers defined recursively: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):1\leq i\leq s\}$, where $1\leq s<r$, $\beta:=\alpha-\sum_{i=1}^{s}q_{i}^{-1}$ belongs to ${\mathscr{W}}_{r}(\alpha)$ as long as $\beta>0$. We first show that there exists a tree $T$ such that its maximum degree $\Delta(T)$ is at most $r$ and its spectrum radius $\lambda(T)$ is equal to $\alpha$ if and only if $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. It follows that the set of spectrum radii of non-trivial trees is exactly the set of positive numbers $\alpha$ such that $\alpha^{-1}\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$. Applying this conclusion, we prove that for any positive integers $r$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
AMS Subject classification: 05C05; 05C20; 05C50
Keywords: tree, maximum degree, characteristic polynomial; eigenvalue; spectrum radius
1 Introduction
All graphs considered in this article are simple and undirected. For a graph $G$, let $V(G)$, $E(G)$ and $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively. For any non-empty subset $S$ of $V(G)$, let $G[S]$ denote the subgraph of $G$ induced by $S$. For any $u\in V(G)$, let $N_{G}(u)$ (or simply $N(u)$) denote the set of neighbours of $u$ in $G$, and let $d_{G}(u)$ or (simply $d(u)$) be the size of $N_{G}(u)$, called the degree of $u$ in $G$. Let $L(G)$ denote the set of vertices $v\in V(G)$ with $d_{G}(v)=1$.
For any integer $k>0$, let $\llbracket k\rrbracket=\{1,2,\dots,k\}$. For a graph $G=(V,E)$, where $V=\{v_{i}:i\in\llbracket k\rrbracket\}$, its adjacency matrix $A(G)$ according to the vertex ordering $v_{1},v_{2},\ldots,v_{n}$ is defined to be the $0-1$ matrix $(a_{i,j})_{n\times n}$, where $a_{i,j}=1$ if and only if $v_{i}v_{j}\in E$. The characteristic polynomial of $G$, denoted by $P(G,\lambda)$, is defined to be the function $det(\lambda I_{n}-A(G))$, where $I_{n}$ is the identical matrix of size $n$. It is known that $P(G,\lambda)$ is independent of the vertex ordering $v_{1},v_{2},\dots,v_{n}$, and the roots of the equation $P(G,\lambda)=0$, called the eigenvalues of $G$, are real numbers (see [6]), and the largest eigenvalue of $G$, denoted by $\lambda(G)$, is called the spectral radius of $G$.
This spectral radius of graphs, which has been studied extensively (see a survey [17] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 8]). In this article, we focus on the study of spectral radii of trees. Let $\Lambda$ denote the set of $\lambda(T)$ for all non-trivial trees $T$ (i.e., trees of order at least $2$). It is known that $q\geq 1$ for each $q\in\Lambda$. The first aim in this article is to provide a characterization of the set $\Lambda$. For any positive number $\alpha$ and positive integer $r$, we define a set ${\mathscr{W}}_{r}(\alpha)$ of positive numbers as follows:
  1. (i).
    $\alpha\in{\mathscr{W}}_{r}(\alpha)$; and
  2. (ii).
    for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s<r$, $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}$ belongs to ${\mathscr{W}}_{r}(\alpha)$ whenever $\beta>0$.
It can be verified that ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha>0$, ${\mathscr{W}}_{r}(1)=\{1\}$ for all $r\geq 2$, ${\mathscr{W}}_{2}(\sqrt{2})=\{\sqrt{2},(\sqrt{2})^{-1}\}$ and ${\mathscr{W}}_{2}(2)=\{\frac{i+1}{i}:i\geq 1\}$. In this article, we apply the set ${\mathscr{W}}_{r}(\alpha)$ to determine if there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ for any given pair of positive numbers $\alpha$ and $r$, where $r$ is an integer.
Theorem 1.1.
For any positive number $\alpha$ and positive integer $r$, there exists a tree $T$ such that $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ if and only if $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
For any tree $T$, $|V(T)|\geq 3$ if and only if $\Delta(T)\geq 2$. In 2003, Stevanović [16] proved that for any tree $T$ with $|V(T)|\geq 3$,
$\frac{1}{4}\lambda(T)^{2}+1<\Delta(T)\leq\lambda(T)^{2}.$ (1)
By the definition of ${\mathscr{W}}_{r}(\alpha)$, we have ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r^{\prime}}(\alpha)$ whenever $r<r^{\prime}$. Thus, a characterization of $\Lambda$ follows from Theorem 1.1 and (1) directly.
Corollary 1.2.
$\Lambda$ is exactly the set of positive numbers $\alpha$ such that $\alpha^{-1}\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)$.
By (1), $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$ is a necessary condition for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$. We wonder if, for any $\alpha\in\Lambda$ and positive integer $r$, this condition is sufficient for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
Conjecture 1.3.
Let $\alpha\in\Lambda\setminus\{1\}$. If $r$ is an integer such that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, then there exists a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
Our second aim in this article is to apply Theorem 1.1 to prove Conjecture 1.3 for the case $\alpha=\sqrt{k}$, where $k\geq 2$ is an integer. Certainly, it includes the case $\alpha=k$ for any integer $k\geq 2$.
Theorem 1.4.
For any positive integers $k$ and $r$, where $r\geq 2$, there exists a tree $T$ with $\lambda(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
The remainder of this article is structured as follows. In Section 2, we examine the set ${\mathscr{W}}_{r}(\alpha)$ and establish a connection between a tree $T$ with $\lambda(T)=\alpha$ and the elements of ${\mathscr{W}}_{r}(\alpha)$. In Section 3, we introduce a family of non-trivial directed trees $T$ with a unique source $z$ and a vertex-weight function $w$, ensuring that $w(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)\setminus\{z\}$, for any given positive numbers $r$ and $\alpha$, where $r$ is an integer. In Sections 4 and 5, we present proofs of Theorems 1.1 and 1.4, respectively. Finally, in Section 6, we discuss open problems arising from our work.
2 The set ${\mathscr{W}}_{r}(\alpha)$
Let $\alpha$ and $r$ be positive, where $r$ is an integer. By definition, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ and $0<q\leq\alpha$ for each $q\in{\mathscr{W}}_{r}(\alpha)$. The following result tells that when $r\leq\alpha^{2}/4+1$, all numbers in ${\mathscr{W}}_{r}(\alpha)$ are at least $\frac{\alpha}{2}$.
Lemma 2.1.
If $12\leq r\leq\alpha^{2}/4+$ 1, then $q>\frac{\alpha}{2}$ for each $q\in{\mathscr{W}}_{r}(\alpha)$.
Proof. Let $\beta:=\frac{\alpha+\sqrt{\alpha^{2}-4r+4}}{2}$. Since $12\leq r\leq\alpha^{2}/4+$ 1, we have $\alpha>\beta\geq\frac{\alpha}{2}$.
We will prove that $q>\beta$ holds for each $q\in{\mathscr{W}}_{r}(\alpha)$. Suppose this conclusion does not hold. By the definition of ${\mathscr{W}}_{r}(\alpha)$ and the fact $\alpha>\beta$, there exists a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $1\leq s<r$, such that $q_{i}>\beta$ for all $i\in\llbracket s\rrbracket$ and
$q:=\alpha-\sum_{i=1}^{s}q_{i}^{-1}\leq\beta.$ (2)
However, it contradicts the following conclusion:
$q=\alpha-\sum_{i=1}^{s}q_{i}^{-1}>\alpha-(r-1)\beta^{-1}=\beta,$ (3)
where the equality $\alpha-(r-1)\beta^{-1}=\beta$ follows from the assumption $\beta=\frac{\alpha+\sqrt{\alpha^{2}-4r+4}}{2}$. Hence the lemma follows. $\Box$
For example, if $r=2$ and $\alpha=2$, then $r=\frac{\alpha^{2}}{4}+1$. By Lemma 2.1, $q>\frac{2}{2}=1$ for all $q\in{\mathscr{W}}_{2}(2)$. This conclusion coincides with the fact that ${\mathscr{W}}_{2}(2)=\{\frac{i+1}{i}:i\geq 1\}$. By Theorem 1.1, Lemma 2.1 implies that there is no trees $T$ with $12\leq\Delta(T)\leq\lambda(T)^{2}/4+$ 1, which coincides with inequality (1) due to Stevanović [16].
Now we are going to introduce the Perron–Frobenius theorem, by which we develop a result for studying the numbers in ${\mathscr{W}}_{r}(\alpha)$.
The Perron–Frobenius theorem, proved by Oskar Perron [14] in 1907 and Georg Frobenius [9] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices. Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectrum radius $\lambda(G)$ of a connected graph $G$ is obtained. Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$. Then, $\lambda(G)=\lambda_{1}(G)$. We say $G$ is non-trivial if $|V(G)|\geq 2$.
Theorem 2.2 (Perron–Frobenius theorem).
Let $G$ be a connected non-trivial graph. Then $\lambda_{1}(G)>\max\{0,\lambda_{2}(G)\}$ and there is an eigenvector $\bar{x}$ of $A(G)$ corresponding to $\lambda_{1}(G)$ in which all components are positive. Furthermore, $\lambda_{1}(G)$ is the only eigenvalue of $G$ with this property.
Applying Theorem 2.2, the next conclusion follows directly. Let ${\mathbb{R}}^{+}$ be the set of positive real numbers.
Proposition 2.3.
Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number. Then, $\lambda(G)=\alpha$ if and only if there is a mapping $\phi$ from $V(G)$ to ${\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
$\alpha\phi(u)=\sum_{w\in N(u)}\phi(w).$ (4)
Now we are going to establish a relation between a mapping $\phi:V(G)\rightarrow{\mathbb{R}}^{+}$ satisfying the condition in (4) and the numbers in ${\mathscr{W}}_{r}(\alpha)$.
Lemma 2.4.
Let $T$ be a non-trivial tree and $\alpha=\lambda(T)$. If $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is a mapping satisfying the condition that for each $u\in V(T)$,
$\alpha\phi(u)=\sum_{w\in N_{T}(u)}\phi(w),$ (5)
then $\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$ for each edge $u_{1}u_{2}$ in $T$, where $r=\Delta(T)$. Therefore, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
Proof. Let $u_{1}u_{2}$ be a fixed edge in $T$ and let $T^{\prime}$ be the subtree of $T$ induced by $\{u_{1}\}\cup V(T_{2})$, where $T_{i}$ is the component of $T-u_{1}u_{2}$ (i.e., the subgraph of $T$ obtained by deleting edge $u_{1}u_{2}$) that contains vertex $u_{i}$. Clearly, $u_{1}\in L(T^{\prime})$ (i.e., a leaf in $T^{\prime}$). For convenience, let $T^{\prime}$ denote the directed tree obtained from $T^{\prime}$ such that $u_{1}$ is its unique source (i.e., a vertex with in-degree 00), as shown in Figure 1 (b).
We are now going to prove the following claim by induction. For any vertex $w$ in $T^{\prime}$, let $d_{T^{\prime}}(u_{1},w)$ denote the distance from $u_{1}$ to $w$ in $T^{\prime}$.
Claim 1: For any directed edge $(w_{1},w_{2})$ in $T^{\prime}$, $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$.
Let $d=\max\{d_{T^{\prime}}(u_{1},w):w\in V(T^{\prime})\}$. If $w_{2}$ is a sink of $T^{\prime}$ (i.e., $w_{2}\in L(T)$), then by condition (5), $\alpha\phi(w_{2})=\phi(w_{1})$, implying that $\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha\in{\mathscr{W}}_{r}(\alpha)$. Thus, Claim 1 holds for any directed edge $(w_{1},w_{2})$ of $T^{\prime}$ with $d_{T^{\prime}}(u_{1},w_{2})=d$.
[…]
    (a) $T$                                                                  (b)     $T^{\prime}$
Figure 1: Tree $T$ and directed tree $T^{\prime}$ in ${\mathscr{T}}$
Assume that Claim 1 holds for each directed edge $(w^{\prime}_{1},w^{\prime}_{2})$ in $T^{\prime}$ with $d_{T^{\prime}}(u_{1},w^{\prime}_{2})\geq t$, where $2\leq t\leq d$. Now we take any directed edge $(w_{1},w_{2})$ in $T^{\prime}$ such that $w_{2}$ is not a sink of $T^{\prime}$ and $d_{T^{\prime}}(u_{1},w_{2})=t-1$. Let $N_{T}(w_{2})\setminus\{w_{1}\}=\{v_{i}:i\in\llbracket s\rrbracket\},$ where $s=d_{T}(w_{2})-1\leq r-1$. Clearly, $d_{T^{\prime}}(u_{1},v_{i})=d_{T^{\prime}}(u_{1},w_{2})+1=t$ for all $i\in\llbracket s\rrbracket$. By the inductive assumption, $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$. By condition (5),
$\alpha\phi(w_{2})=\phi(w_{1})+\sum_{i=1}^{s}\phi(v_{i}),$ (6)
implying that
$\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha-\sum_{i=1}^{s}\frac{\phi(v_{i})}{\phi(w_{2})}=\alpha-\sum_{i=1}^{s}\left(\frac{\phi(w_{2})}{\phi(v_{i})}\right)^{-1}.$ (7)
As $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$, by the definition of ${\mathscr{W}}_{r}(\alpha)$, we have $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Hence Claim 1 holds.
It follows from Claim 1 that $\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$ holds for each edge $u_{1}u_{2}$ in $T$. Taking any edge $u_{1}u_{2}\in E(T)$, where $u_{1}\in L(T)$, we have $\alpha\phi(u_{1})=\phi(u_{2})$ by (5), implying that $\alpha^{-1}=\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Thus, this result follows. $\Box$
3 Weighted Directed trees
Let ${\mathscr{T}}$ be the set of directed trees $T$ of order at least $2$ which has a unique source. Clearly, for any $T\in{\mathscr{T}}$ with the unique source $z$, the in-degree of each vertex $v\in V(T)\setminus\{z\}$ in $T$ is $1$. For any undirected non-trivial tree and any vertex $z$ in this tree, there is one and only one orientation to obtain a directed tree with $z$ as its unique source.
For any real number $\alpha$ with $\alpha\geq 1$, let ${\mathscr{T}}_{\alpha}$ be the set of trees $T\in{\mathscr{T}}$ such that (8) defines a vertex-weight function $\omega$ from $V(T)$ to $\{0\}\cup{\mathbb{R}}^{+}$:
$\omega(v)=\left\{\begin{array}[]{ll}\alpha,&\mbox{if }od_{T}(v)=0;\\ \alpha-\sum\limits_{w\in V(T)\atop(v,w)\in E(T)}\omega(w)^{-1},&\mbox{otherwise},\end{array}\right.$ (8)
where $od_{T}(v)$ is the out-degree of $v$ in $T$ and $(v,w)\in E(T)$ represents the directed edge from $v$ to $w$ in $T$. For any $T\in{\mathscr{T}}_{\alpha}$, let $\omega_{\alpha,T}$ denote the mapping defined in (8), which is unique with respect to $\alpha$ and $T$. Clearly, $\omega_{\alpha,T}(v)>0$ for all $v\in V(T)$, unless $v$ is the unique source of $T$.
For example, if $T\in{\mathscr{T}}$ is the directed tree obtained from the star $K_{1,3}$ so that its unique source is of out-degree $1$ in $T$, it can be verified that $T$ belongs to both ${\mathscr{T}}_{2}$ and ${\mathscr{T}}_{\sqrt{3}}$, as shown in Figure 2 (a) and (b), but $T\notin{\mathscr{T}}_{\frac{3}{2}}\cup{\mathscr{T}}_{\sqrt{2}}$, as shown in Figure 2 (c) and (d), where the number beside each vertex $v$ in $T$ is the value of $\omega_{\alpha,T}(v)$.
1111111122222222
[…]
$-\frac{9}{2}$$\frac{1}{6}$$\frac{3}{2}$$\frac{3}{2}$
[…]
(a) $T\in{\mathscr{T}}_{2}$                  (b) $T\in{\mathscr{T}}_{\sqrt{3}}$                  (c) $T\notin{\mathscr{T}}_{\frac{3}{2}}$                  (d) $T\notin{\mathscr{T}}_{\sqrt{2}}$
Figure 2: $T\in{\mathscr{T}}_{2}$, $T\in{\mathscr{T}}_{\sqrt{3}}$, but $T\notin{\mathscr{T}}_{\frac{3}{2}}\cup{\mathscr{T}}_{\sqrt{2}}$, where $T$ is oriented from $K_{1,3}$
For non-negative numbers $r$, $\alpha$ and $q$, where $r$ is a positive integer and $\alpha\geq 1$, let ${\mathscr{T}}_{r,\alpha}(q)$ denote the set of $T\in{\mathscr{T}}_{\alpha}$ satisfying the following conditions:
  1. (i).
    $\omega_{\alpha,T}(z)=q$, where $z$ is the unique source of $T$, and
  2. (ii).
    for each $v\in V(T)$, $od_{T}(v)\leq r$ if $q=0$ and $v=z$, and $od_{T}(v)\leq r-1$ otherwise.
For example, the directed tree in Figure 2 belongs to both ${\mathscr{T}}_{3,2}(1)$ and ${\mathscr{T}}_{3,\sqrt{3}}(0)$. Two directed trees of ${\mathscr{T}}_{3,2}(\frac{1}{2})$ are also shown in Figure 3.
$\frac{1}{2}$1222
[…]
    (a)                                                                  (b)
Figure 3: Two directed trees in $\mathcal{J}_{3,2}(\frac{1}{2})$
Lemma 3.1.
Let $\alpha\geq 1$, $q\geq 0$ and $r$ be a positive integer. For any $T\in{\mathscr{T}}_{r,\alpha}(q)$, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$, unless $q=0$ and $v$ is the unique source of $T$.
Proof. Since $T\in{\mathscr{T}}_{r,\alpha}(q)$, we have $T\in{\mathscr{T}}_{\alpha}$. In the definition of ${\mathscr{T}}_{\alpha}$, the condition at (8) for the mapping $\omega_{\alpha,T}$ from $V(T)$ to $\{0\}\cup{\mathbb{R}}^{+}$ implies that for any vertex $v$ in $T$, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ holds, unless $v$ is the source of $T$ and $q=0$. Hence the result follows. $\Box$
4 Proof of Theorem 1.1
Lemma 4.1.
Let $\alpha\geq 1$ and $r$ be a positive integer. If there exists a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s\leq r$, such that $\sum\limits_{1\leq i\leq s}q_{i}^{-1}=\alpha$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$.
Proof. By the definition of ${\mathscr{W}}_{r}(\alpha)$ and (8), it is clear that ${\mathscr{T}}_{r,\alpha}(q)\neq\emptyset$ for any $q\in{\mathscr{W}}_{r}(\alpha)$.
Assume that $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$. For each $i\in\llbracket s\rrbracket$, let $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$, $\omega_{i}$ be the mapping $\omega_{\alpha,T_{i}}$ and $z_{i}$ be the unique source of $T_{i}$. By definition, $\omega_{i}(z_{i})=q_{i}$ for all $i\in\llbracket s\rrbracket$.
Now let $T$ be the directed tree obtained from $T_{1},T_{2},\dots,T_{s}$ by adding a new vertex $z$ and adding a new directed edge $(z,z_{i})$ for each $i\in\llbracket s\rrbracket$. Let $\omega:V(T)\mapsto{\mathbb{R}}^{+}\cup\{0\}$ be the mapping defined by $\omega(z)=0$ and $\omega(w)=\omega_{i}(w)$ whenever $w\in V(T_{i})$. Since $\sum\limits_{1\leq i\leq s}q_{i}^{-1}=\alpha$, condition (8) holds for $\omega$ at $z$. For any $v\in V(T)$, if $v\in V(T_{i})$, as $\omega(w)=\omega_{i}(w)$ for all $w\in V(T_{i})$, condition (8) also holds $\omega$ at $v$.
Since $s\leq r$ and $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket s\rrbracket$, we have $\Delta(T)\leq r$, we have $T\in{\mathscr{T}}_{r,\alpha}(0)$. $\Box$
Applying the set ${\mathscr{T}}_{r,\alpha}(0)$, we can now obtain the following conclusion for proving Theorem 1.1.
Lemma 4.2.
Let $\alpha\geq 1$ and $r$ be a positive integer. For any $T\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\lambda(T)=\alpha$. If $\alpha=1$, then $T\in{\mathscr{T}}_{r,\alpha}(0)$ implies that $|V(T)|=2$, and the result holds. Now assume that $\alpha>1$.
Let $T\in{\mathscr{T}}_{r,\alpha}(0)$. By definition, $T$ is a directed tree with a unique source, say $z$. By Lemma 3.1, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)\setminus\{z\}$. Let $T_{1},T_{2},\dots,T_{s}$ be the components of $T-z$. Clearly, $1\leq s\leq r$ and $T_{i}\in{\mathscr{T}}_{\alpha}$ for each $i\in\llbracket s\rrbracket$. Let $z_{i}$ be the vertex in $T_{i}$ such that $(z,z_{i})$ is a directed edge in $T$. Then, $z_{i}$ the unique source of $T_{i}$ and $q_{i}=\omega_{\alpha,T}(z_{i})$ for each $i\in\llbracket s\rrbracket$. By (8), we have
$0=\omega_{\alpha,T}(z)=\alpha-\sum_{i=1}^{s}q_{i}^{-1}.$ (9)
In order to prove that $\lambda(T)=\alpha$, by Proposition 2.3, it suffices to show that the mapping $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ defined below satisfies condition (4):
  1. (i).
    $\phi(z)=1$; and
  2. (ii).
    for each directed edge $(w_{1},w_{2})\in E(T)$, $\phi(w_{2})=\frac{\phi(w_{1})}{\omega(w_{2})}$, where $\omega$ simply denotes $\omega_{\alpha,T}$.
We can first show that condition (4) is satisfied for the unique source and all sinks of $T$. As $z$ is the only source of $T$ and $N^{+}_{T}(z)=\{z_{i}:i\in\llbracket s\rrbracket\}$, where $N^{+}_{T}(z)$ is the set of out-neighbors of $z$ in $T$, we have
$\sum_{v\in N_{T}(z)}\phi(v)=\sum_{i=1}^{s}\phi(z_{i})=\sum_{i=1}^{s}\frac{\phi(z)}{\omega(z_{i})}=\phi(z)\sum_{i=1}^{s}q_{i}^{-1}=\alpha\phi(z),$ (10)
where the last step follows from (9).
For any sink $w$ of $T$, $\omega(w)=\alpha$ by (8). If $w^{\prime}$ is the unique in-neighbor of $w$ in $T$, then $\phi(w^{\prime})=\phi(w)\omega(w)$ by the definition of $\phi$, and
$\sum_{v\in N_{T}(w)}\phi(v)=\phi(w^{\prime})=\phi(w)\omega(w)=\alpha\phi(w).$ (11)
It remains to consider any vertex $w$ in $T$ which is neither the source nor a sink of $T$. Assume that $w^{\prime}$ ie the only in-neighbor of $w$ and $N^{+}_{T}(w)=\{w_{1},w_{2},\dots,w_{t}\}$. By the definition of $\phi$,
$\phi(w^{\prime})+\sum_{i=1}^{t}\phi(w_{i})=\omega(w)\phi(w)+\sum_{i=1}^{t}\frac{\phi(w)}{\omega(w_{i})}=\phi(w)\left(\omega(w)+\sum_{i=1}^{t}\omega(w_{i})^{-1}\right)=\alpha\phi(w),$ (12)
where the last step follows from (8). Thus, $\phi$ satisfies condition (4) for all vertices in $T$, implying that $\lambda(T)=\alpha$ by Proposition 2.3. $\Box$
Now we are going to establish the following conclusion, by which Theorem 1.1 follows directly.
Proposition 4.3.
For any positive numbers $\alpha$ and $r$, where $\alpha\geq 1$ and $r$ is an integer, the following three statements are pairwise equivalent:
  1. (i).
    there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$;
  2. (ii).
    $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$; and
  3. (iii).
    there exist $q_{1},q_{2},\dots,q_{s}\in{\mathscr{W}}_{r}(\alpha)$, where $1\leq s\leq r$, such that $\sum\limits_{i=1}^{s}q_{i}^{-1}=\alpha$.
Proof. (iii) follows from (ii) directly. By Lemmas 4.2 and 4.1, (i) follows from (iii).
(i) $\Rightarrow$ (ii) Assume that $T$ is a tree with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$. By Proposition 2.3, there exists a mapping $\phi$ from $V(T)$ to ${\mathbb{R}}^{+}$ such that $\alpha\phi(u)=\sum\limits_{w\in N_{T}(u)}\phi(w)$ for each $u\in V(T)$. Then, by Lemma 2.4, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. Thus, (i) implies (ii). $\Box$
Clearly, Theorem 1.1 follows from Proposition 4.3 directly. We end this section by a result on the existence of trees $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
For any directed graph $D$, let $\Delta^{(o)}(D)$ denote the maximum out-degree of $D$. By Lemma 4.2 and its proof, the following conclusion holds.
Proposition 4.4.
Let $\alpha\geq 1$ and $r$ be an integer with $\alpha^{2}/4+1<r\leq\alpha^{2}$. Assume that there exist a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s\leq r$, such that $\sum\limits_{i=1}^{s}q_{i}^{-1}=\alpha$. If $s=r$ or $\Delta^{(o)}(T_{i})=r-1$ for some $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$, where $i\in\llbracket s\rrbracket$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
Corollary 4.5.
For any positive real number $\alpha$ and positive integer $r$ with $\alpha^{2}/4+1<r\leq\alpha^{2}$, if there exist a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket r\rrbracket\}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
5 Proof of Theorem 1.4
Let $k$ be an integer with $k\geq 2$ and $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. Clearly, for any integer $r$, $r>\frac{1}{4}k+1$ if and only if $r\geq r_{0}$. If $k=4m+t$, where $0\leq t\leq 3$, then $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2=m+2$.
In order to prove Theorem 1.4, by Corollary 4.5, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{r}(\sqrt{k})$) for all integers $s$ in the interval $[r_{0},k]$. We start with the following simple result.
Lemma 5.1.
Let $k=4m+t\geq 2$, where $m$ and $t$ are integers with $m\geq 0$ and $0\leq t\leq 3$.
  1. (i).
    When $m=0$ and $1\leq t\leq 3$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[1,t]$; and
  2. (ii).
    when $m\geq 1$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[3m+t-1,4m+t]$.
Proof. (i). As $m=0$, we have $k=t$ and thus, $\sqrt{k}=\sqrt{t}$. By definition, $\frac{t}{\sqrt{t}}=\sqrt{t}\in{\mathscr{W}}_{m+2}(\sqrt{t})$ and for $2\leq t\leq 3$,
$\frac{t-1}{\sqrt{t}}=\sqrt{t}-\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t}).$
Thus, it mains to show that $\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t})$ for $t=3$, i.e., $\frac{1}{\sqrt{3}}\in{\mathscr{W}}_{m+2}(\sqrt{3})$. Note that $\frac{2}{\sqrt{3}}\in{\mathscr{W}}_{m+2}(\sqrt{3})$. Thus,
$\frac{\sqrt{3}}{2}=\sqrt{3}-\frac{1}{\frac{2}{\sqrt{3}}}\in{\mathscr{W}}_{m+2}(\sqrt{3}),$
and
$\frac{1}{\sqrt{3}}=\sqrt{3}-\frac{1}{\frac{\sqrt{3}}{2}}\in{\mathscr{W}}_{m+2}(\sqrt{3}).$
(ii) Assume that $m\geq 1$. For any integer $s$ with $3m+t-1\leq s\leq 4m+t$, we have $10\leq 4m+t-s\leq m+$ 1 and
$\frac{s}{\sqrt{4m+t}}=\sqrt{4m+t}-\frac{4m+t-s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{k}).$ (13)
Thus, the result holds. $\Box$
We are now going to establish the conclusion that for any integer $k=4m+t$, where $m\geq 1$ and $t\in\{0,1,2,3\}$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ holds for all integers $s$ in the interval $[m,4m+t]$, on a case-by-case basis of $t$.
Proposition 5.2.
For any positive integer $m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $s$ in $[m,4m]$.
Proof. We will prove that the conclusion in the following claims. This result is true for integers $s$ in the interval $[3m,4m]$ by Lemma 5.1.
Claim 1: $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i\geq 1$.
By definition, $2\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. Thus, the claim holds for $i=1$. For any positive integer $i$, if $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then by definition,
$\frac{i+2}{i+1}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{i+1}{i}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (14)
Thus, the claim holds.
Claim 2: $\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ and $\frac{m-1}{m}{\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
By Claim 1, $\frac{m+1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. Thus,
$\sqrt{m}=2\sqrt{m}-\frac{m+1}{\frac{m+1}{m}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (15)
and
$\frac{m-1}{m}{\sqrt{m}}=2\sqrt{m}-\frac{m+1}{\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (16)
Thus, Claim 2 holds.
Claim 3: $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i$ with $20\leq i\leq m-$ 2.
By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. For any integer $i$ with $21\leq i\leq m-$ 2, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
$\frac{m-i-1}{m-i}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{m-i}{m-i+1}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (17)
Thus, the claim holds.
Claim 4: For any integer $s$ with $m\leq s\leq 4m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
By Lemma 5.1, Claim 4 holds for $3m-1\leq s\leq 4m$. For any integer $i$ with $20\leq i\leq m-$ 2, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
$\frac{2m+i}{2{\sqrt{m}}}=2\sqrt{m}-\frac{m-i-1}{\frac{(m-i-1){\sqrt{m}}}{m-i}}-\frac{i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (18)
Hence Claim 4 holds for $22m\leq s\leq 3m-$ 2. In particular, $\sqrt{m}=\frac{2m}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
For any integer $i$ with $10\leq i\leq m-$ 1, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
$\frac{2{\sqrt{m}}\,(m-i-1)}{2m-i}=2\sqrt{m}-\frac{m+1}{\frac{(2m-i)}{2\sqrt{m}}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ (19)
and
$\frac{2m-i-1}{2{\sqrt{m}}}=2\sqrt{m}-\frac{m-i-1}{\frac{2{\sqrt{m}}\,(m-i-1)}{2m-i}}-\frac{i}{\sqrt{m}}-\frac{1}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (20)
Thus, $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all $i$ with $0\leq i\leq m$. Claim 4 holds.
By Claim 4, the conclusion holds. $\Box$
Proposition 5.3.
For any positive integer $m$, $\frac{s}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for all integers $s$ in $[m,4m+1]$.
Proof. We first prove the following claims.
Claim 1: For all $i\geq 1$, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ and $\frac{(2i+1)m}{i\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
Let $Q(i)=\frac{(2i+1)m}{i\sqrt{4m+1}}$. By Lemma 5.1, $Q(1)=\frac{3m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ By definition, if $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then,
$b:=\sqrt{4m+1}-\frac{m}{Q(i)}=\sqrt{4m+1}-\frac{m}{\frac{(2i+1)m}{i\sqrt{4m+1}}}=\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ (21)
and
$Q(i+1)=\frac{(2i+3)m}{(i+1)\sqrt{4m+1}}=\sqrt{4m+1}-\frac{m}{b}-\frac{1}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (22)
Thus, Claim 1 follows.
Claim 2: For any integer $i$ with $1\leq i\leq m$, we have $\frac{3m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
By Claim 1, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for $i\geq 1$. If $1\leq i\leq m$, then
$\frac{3m-i}{\sqrt{4m+1}}=\sqrt{4m+1}-\frac{i+1}{\frac{(i+1)\sqrt{4m+1}}{2i+1}}-\frac{m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (23)
Claim 2 follows.
Claim 3: For any integer $i$ with $0\leq i\leq m$, we have $\frac{2m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
Let $P(i)=\frac{2m-i}{\sqrt{4m+1}}$. By Claim 2, $P(0)=\frac{2m}{\sqrt{4m+1}}=\frac{3m-m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. By definition,
$\sqrt{4m+1}-\frac{m+1}{\frac{2m}{\sqrt{4m+1}}}=\frac{(m-1)\sqrt{4m+1}}{2m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ (24)
and thus
$P(1)=\frac{2m-1}{\sqrt{4m+1}}=\sqrt{4m+1}-\frac{m-1}{\frac{(m-1)\sqrt{4m+1}}{2m}}-\frac{2}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (25)
Hence the claim holds for $i=0,1$. It suffices to show that for any integer $i$ with $20\leq i\leq m-$ 2, if $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. Assume that $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then
$b:=\sqrt{4m+1}-\frac{m+1}{P(i)}=\sqrt{4m+1}-\frac{m+1}{\frac{2m-i}{\sqrt{4m+1}}}=\frac{(m-i-1)\sqrt{4m+1}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (26)
By Claim 1, $\frac{(i+2)\sqrt{4m+1}}{2i+3}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. Note that
$\sqrt{4m+1}-\frac{m-i-1}{b}-\frac{i+2}{\frac{(i+2)\sqrt{4m+1}}{2i+3}}$ $=$ $\sqrt{4m+1}-\frac{m-i-1}{\frac{(m-i-1)\sqrt{4m+1}}{2m-i}}-\frac{i+2}{\frac{(i+2)\sqrt{4m+1}}{2i+3}}$ (27)
$=$ $\frac{2m-i-2}{\sqrt{4m+1}},$
implying that $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ by definition. Thus, Claim 3 holds.
By Lemma 5.1 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 5.4.
For any positive integer $m$, $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for all integers $s$ in $[m,4m+2]$.
Proof. We first prove the following claims.
Claim 1: For any integer $i\geq 1$, both $\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$ belong to ${\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
Let $P(i)=\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $Q(i)=\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$. By Lemma 5.1, $P(1)=\frac{3m+1}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ and $Q(1)=\frac{3m+2}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Thus, Claim 1 holds for $i=1$.
If $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$, then
$\sqrt{4m+2}-\frac{m+1}{\frac{(2i+1)m+i}{i\sqrt{4m+2}}}=\frac{\sqrt{4m+2}\,m(i+1)}{i(2m+1)+m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (28)
and
$Q(i+1)=\frac{(2(i+1)+1)m+(i+1)+1}{(i+1)\,\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m}{\frac{\sqrt{4m+2}\,m(i+1)}{i(2m+1)+m}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (29)
If $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$, then
$\sqrt{4m+2}-\frac{m}{\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}}=\frac{\sqrt{4m+2}\,(i+1)(m+1)}{i(2m+1)+m+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (30)
and
$P(i+1)=\frac{(2(i+1)+1)m+(i+1)}{(i+1)\,\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{\sqrt{4m+2}\,(i+1)(m+1)}{i(2m+1)+m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (31)
Thus, Claim 1 holds.
Claim 2: $\frac{2m-i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $i$ with $-2\leq i\leq m+1$.
Let $W(i)=\frac{2m-i}{\sqrt{4m+2}}$. By Claim 1, $W(-2)=\frac{(2m+1)m+m}{m\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. By definition,
$W(-1)=\frac{2m+1}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{2m+2}{\sqrt{4m+2}}}=\sqrt{4m+2}-\frac{m+1}{W(-2)}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (32)
$W(0)=\frac{2m}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{2m+1}{\sqrt{4m+2}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (33)
and
$W(1)=\frac{2m-1}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{1}{\frac{2m+1}{\sqrt{4m+2}}}-\frac{m}{\frac{2m}{\sqrt{4m+2}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (34)
Thus, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for each integer $i$ with $-2\leq i\leq 1$.
It suffices to show that for any $i$ with $10\leq i\leq m-$ 1, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ implies that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Observe that for any $i$ with $10\leq i\leq m-$ 1,
$b:=\sqrt{4m+2}-\frac{m+1}{W(i)}=\sqrt{4m+2}-\frac{m+1}{\frac{2m-i}{\sqrt{4m+2}}}=\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (35)
and
$W(i+2)=\frac{2m-i-2}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m-i-1}{b}-\frac{i+2}{W(-1)},$ (36)
implying that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Thus, Claim 2 holds.
Claim 3: $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $s$ with $2m+1\leq s\leq 3m$.
By Claim 2, $\frac{2m+i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for $i=0,1$. For any integer $i$ with $20\leq i\leq m-$ 2, it is shown in (35) that $\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ Thus,
$\frac{2m+2+i}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m-i-1}{\frac{(m-i-1)\cdot\sqrt{4m+2}}{2m-i}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (37)
Hence Claim 3 holds.
By Lemma 5.1 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 5.5.
For any positive integer $m$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $s$ in $[m,4m+3]$.
Proof. We first prove the following claims.
Claim 1: For any positive integer $i$, $\frac{(i+1)\sqrt{4m+3}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ and $\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
Let $f_{1}(i)=\frac{(i+1)\sqrt{4m+3}}{2i+1}$ and $f_{2}(i)=\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}$. Obviously, $\sqrt{4m+3}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$,
$f_{2}(1)=\frac{3(m+1)}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ (38)
and
$f_{1}(1)=\frac{2\sqrt{4m+3}}{3}=\sqrt{4m+3}-\frac{m+1}{\frac{3(m+1)}{\sqrt{4m+3}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (39)
Thus, Claim 1 holds for $i=1$.
Now assume that $t$ is an integer with $t\geq 2$ such that $f_{1}(i),f_{2}(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $i$ with $11\leq i\leq t-$ 1. Then, by definition,
$\frac{2mi+i+m}{i\sqrt{4m+3}}=\sqrt{4m+3}-\frac{i+1}{\frac{(i+1)\sqrt{4m+3}}{2i+1}}-\frac{m-i}{\frac{i\sqrt{4m+3}}{2i-1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (40)
$\sqrt{4m+3}-\frac{m+1}{\frac{2mi+i+m}{i\sqrt{4m+3}}}=\frac{(i+1)m\sqrt{4m+3}}{2mi+i+m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (41)
$f_{2}(i+1)=\frac{(2i+3)(m+1)}{(i+1)\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m}{\frac{(i+1)m\sqrt{4m+3}}{2mi+i+m}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (42)
and
$f_{1}(i+1)=\frac{(i+2)\sqrt{4m+3}}{2i+3}=\sqrt{4m+3}-\frac{m+1}{f_{2}(i+1)}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (43)
Thus, Claim 1 holds.
Claim 2: For any integer $i$ with $-3\leq i\leq m$, $\frac{2m-i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
Let $g(i)=\frac{2m-i}{\sqrt{4m+3}}$. By Claim 1, $g(-3)=\frac{2m+3}{\sqrt{4m+3}}=f_{2}(m+1)$. By definition,
$\sqrt{4m+3}-\frac{m+1}{f_{1}(m)}=\sqrt{4m+3}-\frac{m+1}{\frac{(m+1)\sqrt{4m+3}}{2m+1}}=\frac{2(m+1)}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (44)
$\sqrt{4m+3}-\frac{m+1}{\frac{2(m+1)}{\sqrt{4m+3}}}=\frac{\sqrt{4m+3}}{2}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (45)
$\sqrt{4m+3}-\frac{m+1}{\frac{\sqrt{4m+3}}{2}}=\frac{2m+1}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (46)
$\sqrt{4m+3}-\frac{m+1}{\frac{2m+1}{\sqrt{4m+3}}}=\frac{m\sqrt{4m+3}}{2m+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (47)
and
$\sqrt{4m+3}-\frac{m}{\frac{m\sqrt{4m+3}}{2m+1}}-\frac{1}{\frac{\sqrt{4m+3}}{2}}=\frac{2m}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (48)
implying that $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each $i\in\{-2,-1,0\}$.
It remains to show that for any integer $i$ with $10\leq i\leq m-$ 1, $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ implies that $g(i+1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Observe that
$\sqrt{4m+3}-\frac{m+1}{g(i)}=\sqrt{4m+3}-\frac{(m+1)}{\frac{2m-i}{\sqrt{4m+3}}}=\frac{(m-i-1)\sqrt{4m+3}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (49)
By (45), $\frac{\sqrt{4m+3}}{2}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Thus,
$g(i+1)=\frac{2m-i-1}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m-i-1}{\frac{(m-i-1)\sqrt{4m+3}}{2m-i}}-\frac{i+2}{\frac{\sqrt{4m+3}}{2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (50)
Hence Claim 2 follows.
Claim 3: For any integer $s$ with $32m\leq s\leq 4m+$ 3, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
By Lemma 5.1, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $33m+2\leq s\leq 4m+$ 3.
Let $h(i)=\frac{2m+i}{\sqrt{4m+3}}$. By Claim 2, $h(0),h(1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. For any integer $i$ with $0\leq i\leq m$, by Claim 1, $\frac{(m-i+1)\sqrt{4m+3}}{2(m-i)+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$, implying that
$h(i+2)=\frac{2m+2+i}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{(m+1-i)}{\frac{(m-i+1)\sqrt{4m+3}}{2\cdot(m-i)+1}}-\frac{i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (51)
Thus, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $12m\leq s\leq 3m+$ 1. Hence Claim 3 holds. The result then follows from Claims 2 and 3. $\Box$
We are now going to prove Theorem 1.4.
Proof of Theorem 1.4: The necessity follows from (1). Now we prove the sufficiency.
Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. By Lemma 5.1 and Propositions 5.25.35.4 and 5.5, for each integer $s$ with $\left\lfloor{\frac{1}{4}k}\right\rfloor\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ holds, implying that
$s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}.$ (52)
Then, by Corollary 4.5, for any integer $r$ in $[r_{0},k]$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$. $\Box$
6 Problems and Remarks
Recall that $\Lambda$ is the set of $\lambda(T)$ over all trees $T$ of order at least $2$. Conjecture 1.3 asks an interesting question on the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$ for any given $\alpha\in\Lambda\setminus\{1\}$ and integer $r$ satisfying $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
Another interesting problem is on the numbers contained in $\Lambda$. Let ${\cal P}$ be the set of real-rooted monic polynomials with integer coefficients. Salez [15] showed that every totally real algebraic integer (i.e. a root of some polynomial in ${\cal P}$) is an eigenvalue of some finite tree. Now let $\Lambda^{\prime}$ be the set of real numbers $\alpha$ such that $\alpha$ is the largest real root of some $P\in{\cal P}$.
Problem 6.1.
Is there a positive real number $c$ such that $\{\alpha\in\Lambda^{\prime}:\alpha\geq c\}\subseteq\Lambda$?
It can be proved by applying Proposition 4.4 that, for any positive integers $p$ and $q$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if $4|(p-q)$, $\sqrt{p+\sqrt{q}}\in\Lambda$ if $p\geq q$, and $\frac{\sqrt{2(2p+1)+2\sqrt{4q+1}}}{2}\in\Lambda$ if $q\leq p^{2}+p$. As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ whenever $4|(p-q)$.
Proposition 6.2.
For any positive integers $p$ and $q$ with $p\geq q$, if $p-q$ is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\{q,\frac{p-q}{4}\}$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
Proof. Assume that $p-q=4d$. Let $r=\max\{q,d\}+1$ and $\alpha=\frac{\sqrt{p}+\sqrt{q}}{2}$. By definition,
$\sqrt{q}=\frac{\sqrt{p}+\sqrt{q}}{2}-\frac{d}{\frac{\sqrt{p}+\sqrt{q}}{2}}\in{\mathscr{W}}_{r}(\alpha)$ (53)
and
$\frac{\sqrt{p}-\sqrt{q}}{2}=\frac{\sqrt{p}+\sqrt{q}}{2}-\frac{q}{\sqrt{q}}\in{\mathscr{W}}_{r}(\alpha).$ (54)
Observe that
$d\cdot\left(\frac{\sqrt{p}-\sqrt{q}}{2}\right)^{-1}=\frac{\frac{p-q}{4}}{\frac{\sqrt{p}-\sqrt{q}}{2}}=\frac{\sqrt{p}+\sqrt{q}}{2}=\alpha.$ (55)
By Proposition 4.4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
We wonder if $4|(p-q)$ is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\Lambda$. For any positive integers $s$ and $t$, $\frac{s+t}{2}=\frac{\sqrt{s^{2}}+\sqrt{t^{2}}}{2}$. As $P(G,\lambda)\in{\cal P}$, each rational root of $P(G,\lambda)=0$ is an integer. It follows that $\frac{s+t}{2}\in\Lambda$ if and only if $s$ and $t$ have the same parity, i.e., $4|(s^{2}-t^{2})$.
Problem 6.3.
Is it true that for any positive integers $p$ and $q$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if and only if $p-q$ is a multiple of $4$?
Let $T$ be any tree with $\alpha=\lambda(T)$ and $r=\Delta(T)$. Then, (1) yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$. By Theorem 1.1, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
Problem 6.4.
Let $\lambda\in\Lambda$ and $r$ be an integer with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?
Acknowledgement
This research is supported by NSFC (No. 12101347 and 12371340), NSF of Shandong Province (No. ZR2021QA085) and the Ministry of Education, Singapore, under its Academic Research Tier 1 (RG19/22). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore.

### v2

@@
-Spectrum radii of trees
+An algebraic characterization for spectral radii of trees
 Abstract
-For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of positive numbers defined recursively: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):1\leq i\leq s\}$, where $1\leq s<r$, $\beta:=\alpha-\sum_{i=1}^{s}q_{i}^{-1}$ belongs to ${\mathscr{W}}_{r}(\alpha)$ as long as $\beta>0$.
-We first show that there exists a tree $T$ such that its maximum degree $\Delta(T)$ is at most $r$ and its spectrum radius $\lambda(T)$ is equal to $\alpha$ if and only if $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
-It follows that the set of spectrum radii of non-trivial trees is exactly the set of positive numbers $\alpha$ such that $\alpha^{-1}\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$.
-Applying this conclusion, we prove that for any positive integers $r$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
+There have been various results on spectral radii of trees, studying the relations between spectral radius $\lambda(T)$ of a tree $T$ and some other parameters of $T$.
+In this article, we provide an algebraic characterization for the spectral radii of trees, and apply this new result to study the existence of trees with given spectral radius and maximum degree.
+For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set defined recursively: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,1\leq i\leq s\}$, if $\beta:=\alpha-\sum\limits_{i=1}^{s}q_{i}^{-1}\geq 0$ and $s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$.
+We first show that for any positive integer $r$ and real number $\alpha\geq 1$, there exists a tree $T$ with maximum degree $\Delta(T)\leq r$ and spectral radius $\lambda(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
+It follows that the set of spectral radii of non-trivial trees is exactly the set of positive numbers $\alpha$ such that $0\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$.
+Applying this conclusion, we then prove that for any positive integers $r\geq 2$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
 AMS Subject classification: 05C05; 05C20; 05C50
-Keywords: tree, maximum degree, characteristic polynomial; eigenvalue; spectrum radius
+Keywords: tree, maximum degree, characteristic polynomial; eigenvalue; spectral radius
 1 Introduction
@@
 For a graph $G$, let $V(G)$, $E(G)$ and $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively.
-For any non-empty subset $S$ of $V(G)$, let $G[S]$ denote the subgraph of $G$ induced by $S$.
 For any $u\in V(G)$, let $N_{G}(u)$ (or simply $N(u)$) denote the set of neighbours of $u$ in $G$, and let $d_{G}(u)$ or (simply $d(u)$) be the size of $N_{G}(u)$, called the degree of $u$ in $G$.
 Let $L(G)$ denote the set of vertices $v\in V(G)$ with $d_{G}(v)=1$.
-For any integer $k>0$, let $\llbracket k\rrbracket=\{1,2,\dots,k\}$.
+Let ${\mathbb{N}}^{+}$ denote the set of positive integers and for any $k\in{\mathbb{N}}^{+}$, let $\llbracket k\rrbracket=\{i\in{\mathbb{N}}^{+}:i\leq k\}$.
 For a graph $G=(V,E)$, where $V=\{v_{i}:i\in\llbracket k\rrbracket\}$, its adjacency matrix $A(G)$ according to the vertex ordering $v_{1},v_{2},\ldots,v_{n}$ is defined to be the $0-1$ matrix $(a_{i,j})_{n\times n}$, where $a_{i,j}=1$ if and only if $v_{i}v_{j}\in E$.
@@
 It is known that $P(G,\lambda)$ is independent of the vertex ordering $v_{1},v_{2},\dots,v_{n}$, and the roots of the equation $P(G,\lambda)=0$, called the eigenvalues of $G$, are real numbers (see [6]), and the largest eigenvalue of $G$, denoted by $\lambda(G)$, is called the spectral radius of $G$.
-This spectral radius of graphs, which has been studied extensively (see a survey [17] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 8]).
+The spectral radii of graphs, which have been studied extensively (see a survey [26] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 9, 12, 13, 22, 23, 26]).
+The distribution of spectral radii among different graphs is a topic of ongoing research with connections to various graph properties and invariants (see [7, 12, 13, 16, 20, 22, 23]).
 In this article, we focus on the study of spectral radii of trees.
+There are various known results on the relation between $\lambda(T)$ and other parameters of a tree $T$.
+Guo and Tan [11] found an upper bound of $\lambda(T)$ in terms of the order and matching number of $T$.
+Lovász and Pelikán [15] determined the maximal and minimal value of $\lambda(T)$ among all trees $T$ of order $n$, Simić and Tos̆ić [24] found the tree that has the largest spectral radius among all trees with given order and maximum degree, and Stevanović [25] provided a relation between $\lambda(T)$ and $\Delta(T)$: for any tree $T$ with $|V(T)|\geq 3$,
+$\frac{1}{4}\lambda(T)^{2}+1<\Delta(T)\leq\lambda(T)^{2}.$ (1)
+In this article, we will first provide a characterization on $r$ and $\alpha$ for the existence of trees $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
 Let $\Lambda$ denote the set of $\lambda(T)$ for all non-trivial trees $T$ (i.e., trees of order at least $2$).
-It is known that $q\geq 1$ for each $q\in\Lambda$.
-The first aim in this article is to provide a characterization of the set $\Lambda$.
-For any positive number $\alpha$ and positive integer $r$, we define a set ${\mathscr{W}}_{r}(\alpha)$ of positive numbers as follows:
+Clearly, $\alpha\geq 1$ for each $\alpha\in\Lambda$.
+Let ${\mathbb{R}}^{+}$ denote the set of positive real numbers, and for any $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of non-negative real numbers defined below:
   1. (i).
@@
   2. (ii).
-    for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s<r$, $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}$ belongs to ${\mathscr{W}}_{r}(\alpha)$ whenever $\beta>0$.
-It can be verified that ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha>0$, ${\mathscr{W}}_{r}(1)=\{1\}$ for all $r\geq 2$, ${\mathscr{W}}_{2}(\sqrt{2})=\{\sqrt{2},(\sqrt{2})^{-1}\}$ and ${\mathscr{W}}_{2}(2)=\{\frac{i+1}{i}:i\geq 1\}$.
-In this article, we apply the set ${\mathscr{W}}_{r}(\alpha)$ to determine if there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ for any given pair of positive numbers $\alpha$ and $r$, where $r$ is an integer.
+    for any $s\in{\mathbb{N}}^{+}$ and multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket s\rrbracket\big{\}}$, if $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}\geq 0$ and $1\leq s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$.
+It can be verified that ${\mathscr{W}}_{1}(1)=\{0,1\}$, ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha\in{\mathbb{R}}^{+}$ with $\alpha\neq 1$, ${\mathscr{W}}_{2}(\sqrt{2})=\left\{0,\sqrt{2},(\sqrt{2})^{-1}\right\}$ and ${\mathscr{W}}_{2}(2)=\big{\{}\frac{i+1}{i}:i\geq 1\big{\}}$.
+Clearly, the property ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ holds for any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$.
+Our first aim in this article is to establish a connection between the existence of a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ and the existence of 00 in the set ${\mathscr{W}}_{r}(\alpha)$.
 Theorem 1.1.
-For any positive number $\alpha$ and positive integer $r$, there exists a tree $T$ such that $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ if and only if $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
-For any tree $T$, $|V(T)|\geq 3$ if and only if $\Delta(T)\geq 2$. In 2003, Stevanović [16] proved that for any tree $T$ with $|V(T)|\geq 3$,
-$\frac{1}{4}\lambda(T)^{2}+1<\Delta(T)\leq\lambda(T)^{2}.$ (1)
-By the definition of ${\mathscr{W}}_{r}(\alpha)$, we have ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r^{\prime}}(\alpha)$ whenever $r<r^{\prime}$.
-Thus, a characterization of $\Lambda$ follows from Theorem 1.1 and (1) directly.
+For any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$, there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
+A characterization of the set $\Lambda$ follows from Theorem 1.1 and (1) directly.
 Corollary 1.2.
-$\Lambda$ is exactly the set of positive numbers $\alpha$ such that $\alpha^{-1}\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)$.
-By (1), $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$ is a necessary condition for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
-We wonder if, for any $\alpha\in\Lambda$ and positive integer $r$, this condition is sufficient for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
+$\Lambda=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)\big{\}}$.
+If $r$ is removed from the definition of the set ${\mathscr{W}}_{r}(\alpha)$, we can define the set ${\mathscr{W}}(\alpha)$ as follows:
+  1. (i).
+    $\alpha\in{\mathscr{W}}(\alpha)$; and
+  2. (ii).
+    for any multi-set $\big{\{}q_{i}\in{\mathscr{W}}(\alpha):q_{i}>0,i\in\llbracket s\rrbracket\big{\}}$, if $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}\geq 0$, then $\beta\in{\mathscr{W}}(\alpha)$.
+Clearly, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}(\alpha)$ for any $r\in{\mathbb{N}}^{+}$, and the following conclusion holds.
+Corollary 1.3.
+$\Lambda=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}(\alpha)\big{\}}$.
+The conclusion of (1) tells that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$ is a necessary condition for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
+We wonder if this condition is also sufficient.
 Conjecture 1.4.
 Let $\alpha\in\Lambda\setminus\{1\}$. If $r$ is an integer such that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, then there exists a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
-Our second aim in this article is to apply Theorem 1.1 to prove Conjecture 1.3 for the case $\alpha=\sqrt{k}$, where $k\geq 2$ is an integer.
-Certainly, it includes the case $\alpha=k$ for any integer $k\geq 2$.
+Our second aim is to apply Theorem 1.1 to prove Conjecture 1.4 for the case $\alpha=\sqrt{k}$ for each $k\in{\mathbb{N}}^{+}\setminus\{1\}$.
+Certainly, it includes the case $\alpha=k$ for each $k\in{\mathbb{N}}^{+}\setminus\{1\}$.
 Theorem 1.5.
-For any positive integers $k$ and $r$, where $r\geq 2$, there exists a tree $T$ with $\lambda(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
+For any $k,r\in{\mathbb{N}}^{+}\setminus\{1\}$, there exists a tree $T$ with $\lambda(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
 The remainder of this article is structured as follows.
-In Section 2, we examine the set ${\mathscr{W}}_{r}(\alpha)$ and establish a connection between a tree $T$ with $\lambda(T)=\alpha$ and the elements of ${\mathscr{W}}_{r}(\alpha)$.
-In Section 3, we introduce a family of non-trivial directed trees $T$ with a unique source $z$ and a vertex-weight function $w$, ensuring that $w(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)\setminus\{z\}$, for any given positive numbers $r$ and $\alpha$, where $r$ is an integer.
-In Sections 4 and 5, we present proofs of Theorems 1.1 and 1.4, respectively.
-Finally, in Section 6, we discuss open problems arising from our work.
-2 The set ${\mathscr{W}}_{r}(\alpha)$
-Let $\alpha$ and $r$ be positive, where $r$ is an integer.
-By definition, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ and $0<q\leq\alpha$ for each $q\in{\mathscr{W}}_{r}(\alpha)$.
-The following result tells that when $r\leq\alpha^{2}/4+1$, all numbers in ${\mathscr{W}}_{r}(\alpha)$ are at least $\frac{\alpha}{2}$.
-Lemma 2.1.
-If $12\leq r\leq\alpha^{2}/4+$ 1, then $q>\frac{\alpha}{2}$ for each $q\in{\mathscr{W}}_{r}(\alpha)$.
-Proof. Let $\beta:=\frac{\alpha+\sqrt{\alpha^{2}-4r+4}}{2}$. Since $12\leq r\leq\alpha^{2}/4+$ 1, we have $\alpha>\beta\geq\frac{\alpha}{2}$.
-We will prove that $q>\beta$ holds for each $q\in{\mathscr{W}}_{r}(\alpha)$.
-Suppose this conclusion does not hold.
-By the definition of ${\mathscr{W}}_{r}(\alpha)$ and the fact $\alpha>\beta$, there exists a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $1\leq s<r$, such that $q_{i}>\beta$ for all $i\in\llbracket s\rrbracket$ and
-$q:=\alpha-\sum_{i=1}^{s}q_{i}^{-1}\leq\beta.$ (2)
-However, it contradicts the following conclusion:
-$q=\alpha-\sum_{i=1}^{s}q_{i}^{-1}>\alpha-(r-1)\beta^{-1}=\beta,$ (3)
-where the equality $\alpha-(r-1)\beta^{-1}=\beta$ follows from the assumption $\beta=\frac{\alpha+\sqrt{\alpha^{2}-4r+4}}{2}$. Hence the lemma follows. $\Box$
-For example, if $r=2$ and $\alpha=2$, then $r=\frac{\alpha^{2}}{4}+1$.
-By Lemma 2.1, $q>\frac{2}{2}=1$ for all $q\in{\mathscr{W}}_{2}(2)$.
-This conclusion coincides with the fact that ${\mathscr{W}}_{2}(2)=\{\frac{i+1}{i}:i\geq 1\}$.
-By Theorem 1.1, Lemma 2.1 implies that there is no trees $T$ with $12\leq\Delta(T)\leq\lambda(T)^{2}/4+$ 1, which coincides with inequality (1) due to Stevanović [16].
-Now we are going to introduce the Perron–Frobenius theorem, by which we develop a result for studying the numbers in ${\mathscr{W}}_{r}(\alpha)$.
+In Sections 2 and 3, we prove Theorem 1.1’s necessity and sufficiency, respectively.
+In Section 4, we complete the proof of Theorem 1.5.
+Finally, in Section 5, we provide some open problems arising from our work.
+2 The necessity of Theorem 1.1
 The Perron–Frobenius theorem, proved by Oskar Perron [19] in 1907 and Georg Frobenius [10] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices.
-Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectrum radius $\lambda(G)$ of a connected graph $G$ is obtained.
+Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectral radius $\lambda(G)$ of a connected graph $G$ is obtained.
 Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$.
@@
 Furthermore, $\lambda_{1}(G)$ is the only eigenvalue of $G$ with this property.
-Applying Theorem 2.2, the next conclusion follows directly. Let ${\mathbb{R}}^{+}$ be the set of positive real numbers.
+Applying Theorem 2.1, the next conclusion follows directly.
 Proposition 2.2.
 Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number.
-Then, $\lambda(G)=\alpha$ if and only if there is a mapping $\phi$ from $V(G)$ to ${\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
+Then, $\lambda(G)=\alpha$ if and only if there is a mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
 $\alpha\phi(u)=\sum_{w\in N(u)}\phi(w).$ (2)
-Now we are going to establish a relation between a mapping $\phi:V(G)\rightarrow{\mathbb{R}}^{+}$ satisfying the condition in (4) and the numbers in ${\mathscr{W}}_{r}(\alpha)$.
+Note that for any mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$, the condition of (2) holds if and only if $\phi$ is an eigenvector of $A(G)$ with respect to $\lambda(G)$.
+Now we are going to prove the following lemma by which the necessity of Theorem 1.1 follows directly.
 Lemma 2.3.
-Let $T$ be a non-trivial tree and $\alpha=\lambda(T)$. If $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is a mapping satisfying the condition that for each $u\in V(T)$,
-$\alpha\phi(u)=\sum_{w\in N_{T}(u)}\phi(w),$ (5)
-then $\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$ for each edge $u_{1}u_{2}$ in $T$, where $r=\Delta(T)$. Therefore, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
-Proof.
-Let $u_{1}u_{2}$ be a fixed edge in $T$ and let $T^{\prime}$ be the subtree of $T$ induced by $\{u_{1}\}\cup V(T_{2})$, where $T_{i}$ is the component of $T-u_{1}u_{2}$ (i.e., the subgraph of $T$ obtained by deleting edge $u_{1}u_{2}$) that contains vertex $u_{i}$.
-Clearly, $u_{1}\in L(T^{\prime})$ (i.e., a leaf in $T^{\prime}$).
-For convenience, let $T^{\prime}$ denote the directed tree obtained from $T^{\prime}$ such that $u_{1}$ is its unique source (i.e., a vertex with in-degree 00), as shown in Figure 1 (b).
-We are now going to prove the following claim by induction. For any vertex $w$ in $T^{\prime}$, let $d_{T^{\prime}}(u_{1},w)$ denote the distance from $u_{1}$ to $w$ in $T^{\prime}$.
-Claim 1: For any directed edge $(w_{1},w_{2})$ in $T^{\prime}$, $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$.
-Let $d=\max\{d_{T^{\prime}}(u_{1},w):w\in V(T^{\prime})\}$.
-If $w_{2}$ is a sink of $T^{\prime}$ (i.e., $w_{2}\in L(T)$), then by condition (5), $\alpha\phi(w_{2})=\phi(w_{1})$, implying that $\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha\in{\mathscr{W}}_{r}(\alpha)$.
-Thus, Claim 1 holds for any directed edge $(w_{1},w_{2})$ of $T^{\prime}$ with $d_{T^{\prime}}(u_{1},w_{2})=d$.
+Let $T$ be a non-trivial tree and $\alpha=\lambda(T)$, and let $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ be an eigenvector of $A(T)$ with respect to $\alpha$.
+Then both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$, where $r=\Delta(T)$.
+Proof. Since $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is an eigenvector of $A(T)$ with respect to $\alpha=\lambda(T)$, (2) holds for each $u\in V(T)$. Let’s first prove the following claim.
+Claim 1: For an arbitrary vertex $z$ in $T$, if $T^{\prime}$ is the orientation of $T$ with $z$ as its source, as shown in Figure 1, then $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$ holds for every directed edge $(w_{1},w_{2})$ in $T^{\prime}$.
 […]
-    (a) $T$                                                                  (b)     $T^{\prime}$
-Figure 1: Tree $T$ and directed tree $T^{\prime}$ in ${\mathscr{T}}$
-Assume that Claim 1 holds for each directed edge $(w^{\prime}_{1},w^{\prime}_{2})$ in $T^{\prime}$ with $d_{T^{\prime}}(u_{1},w^{\prime}_{2})\geq t$, where $2\leq t\leq d$.
-Now we take any directed edge $(w_{1},w_{2})$ in $T^{\prime}$ such that $w_{2}$ is not a sink of $T^{\prime}$ and $d_{T^{\prime}}(u_{1},w_{2})=t-1$.
-Let $N_{T}(w_{2})\setminus\{w_{1}\}=\{v_{i}:i\in\llbracket s\rrbracket\},$ where $s=d_{T}(w_{2})-1\leq r-1$.
-Clearly, $d_{T^{\prime}}(u_{1},v_{i})=d_{T^{\prime}}(u_{1},w_{2})+1=t$ for all $i\in\llbracket s\rrbracket$.
+Figure 1: A directed tree $T^{\prime}$ with $z$ as its unique source
+We will prove Claim 1 by induction.
+For any vertex $w$ in $T^{\prime}$, let $d_{T^{\prime}}(z,w)$ denote the distance from $z$ to $w$ in $T^{\prime}$, and let $d=\max\{d_{T^{\prime}}(z,w):w\in V(T^{\prime})\}$.
+If $w_{2}$ is a sink of $T^{\prime}$, then $w_{2}$ is a leaf of $T$, and the condition of (2) yields that $\alpha\phi(w_{2})=\phi(w_{1})$, implying that $\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha\in{\mathscr{W}}_{r}(\alpha)$.
+Thus, Claim 1 holds whenever $w_{2}$ is a sink of $T^{\prime}$.
+Clearly, if $d_{T^{\prime}}(u_{1},w_{2})=d$, then $w_{2}$ is a sink of $T^{\prime}$ and thus Claim 1 also holds.
+Assume that Claim 1 holds for each directed edge $(w^{\prime}_{1},w^{\prime}_{2})$ in $T^{\prime}$ with $d_{T^{\prime}}(z,w^{\prime}_{2})\geq t$, where $2\leq t\leq d$.
+Now we take any directed edge $(w_{1},w_{2})$ in $T^{\prime}$ such that $w_{2}$ is not a sink of $T^{\prime}$ and $d_{T^{\prime}}(z,w_{2})=t-1$.
+Assume that $N_{T}(w_{2})\setminus\{w_{1}\}=\{v_{i}:i\in\llbracket s\rrbracket\},$ where $s=d_{T}(w_{2})-1\leq r-1$.
+Clearly, for each $i\in\llbracket s\rrbracket$, $(w_{2},v_{i})$ is a directed edge in $T^{\prime}$ and $d_{T^{\prime}}(z,v_{i})=d_{T^{\prime}}(z,w_{2})+1=t$.
 By the inductive assumption, $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$.
-By condition (5),
+By the condition of (2),
 $\alpha\phi(w_{2})=\phi(w_{1})+\sum_{i=1}^{s}\phi(v_{i}),$ (3)
@@
 $\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha-\sum_{i=1}^{s}\frac{\phi(v_{i})}{\phi(w_{2})}=\alpha-\sum_{i=1}^{s}\left(\frac{\phi(w_{2})}{\phi(v_{i})}\right)^{-1}.$ (4)
-As $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$, by the definition of ${\mathscr{W}}_{r}(\alpha)$, we have $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$.
-Hence Claim 1 holds.
-It follows from Claim 1 that $\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$ holds for each edge $u_{1}u_{2}$ in $T$.
-Taking any edge $u_{1}u_{2}\in E(T)$, where $u_{1}\in L(T)$, we have $\alpha\phi(u_{1})=\phi(u_{2})$ by (5), implying that $\alpha^{-1}=\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$.
-Thus, this result follows.
+As $s<r$, by the definition of ${\mathscr{W}}_{r}(\alpha)$, we have $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Hence Claim 1 holds.
+Since vertex $z$ in Claim 1 is an arbitrary vertex in $T$, Claim 1 implies that both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$.
 $\Box$
-3 Weighted Directed trees
-Let ${\mathscr{T}}$ be the set of directed trees $T$ of order at least $2$ which has a unique source.
-Clearly, for any $T\in{\mathscr{T}}$ with the unique source $z$, the in-degree of each vertex $v\in V(T)\setminus\{z\}$ in $T$ is $1$.
-For any undirected non-trivial tree and any vertex $z$ in this tree, there is one and only one orientation to obtain a directed tree with $z$ as its unique source.
-For any real number $\alpha$ with $\alpha\geq 1$, let ${\mathscr{T}}_{\alpha}$ be the set of trees $T\in{\mathscr{T}}$ such that (8) defines a vertex-weight function $\omega$ from $V(T)$ to $\{0\}\cup{\mathbb{R}}^{+}$:
-$\omega(v)=\left\{\begin{array}[]{ll}\alpha,&\mbox{if }od_{T}(v)=0;\\ \alpha-\sum\limits_{w\in V(T)\atop(v,w)\in E(T)}\omega(w)^{-1},&\mbox{otherwise},\end{array}\right.$ (8)
-where $od_{T}(v)$ is the out-degree of $v$ in $T$ and $(v,w)\in E(T)$ represents the directed edge from $v$ to $w$ in $T$.
-For any $T\in{\mathscr{T}}_{\alpha}$, let $\omega_{\alpha,T}$ denote the mapping defined in (8), which is unique with respect to $\alpha$ and $T$.
-Clearly, $\omega_{\alpha,T}(v)>0$ for all $v\in V(T)$, unless $v$ is the unique source of $T$.
-For example, if $T\in{\mathscr{T}}$ is the directed tree obtained from the star $K_{1,3}$ so that its unique source is of out-degree $1$ in $T$, it can be verified that $T$ belongs to both ${\mathscr{T}}_{2}$ and ${\mathscr{T}}_{\sqrt{3}}$, as shown in Figure 2 (a) and (b), but $T\notin{\mathscr{T}}_{\frac{3}{2}}\cup{\mathscr{T}}_{\sqrt{2}}$, as shown in Figure 2 (c) and (d), where the number beside each vertex $v$ in $T$ is the value of $\omega_{\alpha,T}(v)$.
-1111111122222222
-[…]
-$-\frac{9}{2}$$\frac{1}{6}$$\frac{3}{2}$$\frac{3}{2}$
-[…]
-(a) $T\in{\mathscr{T}}_{2}$                  (b) $T\in{\mathscr{T}}_{\sqrt{3}}$                  (c) $T\notin{\mathscr{T}}_{\frac{3}{2}}$                  (d) $T\notin{\mathscr{T}}_{\sqrt{2}}$
-Figure 2: $T\in{\mathscr{T}}_{2}$, $T\in{\mathscr{T}}_{\sqrt{3}}$, but $T\notin{\mathscr{T}}_{\frac{3}{2}}\cup{\mathscr{T}}_{\sqrt{2}}$, where $T$ is oriented from $K_{1,3}$
-For non-negative numbers $r$, $\alpha$ and $q$, where $r$ is a positive integer and $\alpha\geq 1$, let ${\mathscr{T}}_{r,\alpha}(q)$ denote the set of $T\in{\mathscr{T}}_{\alpha}$ satisfying the following conditions:
+Proof of the necessity of Theorem 1.1.
+We follow the conditions in Lemma 2.3.
+Take any edge $u_{1}u_{2}\in E(T)$, where $u_{1}$ is a leaf of $T$.
+We have $\alpha\phi(u_{1})=\phi(u_{2})$ by (2), implying that $\alpha^{-1}=\frac{\phi(u_{2})}{\phi(u_{1})}\in{\mathscr{W}}_{r}(\alpha)$ by Lemma 2.3.
+Thus, $0=\alpha-\frac{1}{\alpha^{-1}}\in{\mathscr{W}}_{r}(\alpha)$ by definition.
+Hence the result holds. $\Box$
+3 The sufficiency of Theorem 1.1
+In this section, we always assume that $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$.
+Let ${\mathscr{T}}_{r,\alpha}$ denote the set of ordered pairs $(T,\omega)$, where $T$ is a directed tree with a unique source $z$ and $\omega$ is a mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ such that for each $v\in V(T)$,
   1. (i).
-    $\omega_{\alpha,T}(z)=q$, where $z$ is the unique source of $T$, and
+    $od_{T}(v)\leq r$ if $v=z$ and $\omega(z)=0$, and $od_{T}(v)\leq r-1$ otherwise; and
   2. (ii).
-    for each $v\in V(T)$, $od_{T}(v)\leq r$ if $q=0$ and $v=z$, and $od_{T}(v)\leq r-1$ otherwise.
-For example, the directed tree in Figure 2 belongs to both ${\mathscr{T}}_{3,2}(1)$ and ${\mathscr{T}}_{3,\sqrt{3}}(0)$.
-Two directed trees of ${\mathscr{T}}_{3,2}(\frac{1}{2})$ are also shown in Figure 3.
+    $\omega(v)>0$ whenever $v\in V(T)\setminus\{z\}$, and
+    $\omega(v)=\left\{\begin{array}[]{ll}\alpha,&\mbox{if }N^{+}_{T}(v)=\emptyset;\\ \alpha-\sum\limits_{u\in N^{+}_{T}(v)}\omega(u)^{-1},&\mbox{otherwise},\end{array}\right.$ (5)
+where $N^{+}_{T}(v)$ is the set of out-neighbors of $v$ in $T$ and $od_{T}(v)=|N^{+}_{T}(v)|$.
+For any $\beta\in{\mathbb{R}}^{+}$, let ${\mathscr{T}}_{r,\alpha}(\beta)$ denote the set of ordered pairs $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}$ such that $\omega(z)=\beta$, where $z$ is the unique source of $T$.
+By the definitions of ${\mathscr{W}}_{r}(\alpha)$ and ${\mathscr{T}}_{r,\alpha}(\beta)$, the following conclusion holds.
+Lemma 3.1.
+For any $\beta\in{\mathscr{W}}_{r}(\alpha)$, ${\mathscr{T}}_{r,\alpha}(\beta)\neq\emptyset$.
+Proof.
+The conclusion is trivial if $\beta=\alpha$.
+Now assume that $\beta=\alpha-\sum\limits_{i=1}^{s}\beta_{i}^{-1}$ and $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(\beta_{i})$ for each $i\in\llbracket s\rrbracket$, where $s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$.
+Let $T$ be the directed tree obtained from $T_{1},\cdots,T_{s}$ by adding a new vertex $z$ and a new directed edge $(z,z_{i})$ for each $i\in\llbracket s\rrbracket$, where $z_{i}$ is the unique source of $T_{i}$, and let $\omega$ be the mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ such that $\omega(v)=\omega_{i}(v)$ holds for each $i\in\llbracket s\rrbracket$ and $v\in V(T_{i})$, and
+$\omega(z)=\alpha-\sum\limits_{i=1}^{s}\omega(z_{i})^{-1}=\alpha-\sum\limits_{i=1}^{s}\beta_{i}^{-1}=\beta.$ (6)
+Obviously, $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(\beta)$. Hence the lemma holds. $\Box$
+For example, two members $(T_{i},\omega_{i})$ for $i=1,2$ of $\mathcal{J}_{3,2}(\frac{1}{2})$ are shown in Figure 2, where the value of $\omega_{i}(v)$ for each vertex $v$ in $T_{i}$ is shown beside vertex $v$.
 $\frac{1}{2}$1222
 […]
-    (a)                                                                  (b)
-Figure 3: Two directed trees in $\mathcal{J}_{3,2}(\frac{1}{2})$
+    (a) $(T_{1},\omega_{1})$                                             (b) $(T_{2},\omega_{2})$
+Figure 2: $(T_{i},\omega_{i})\in\mathcal{J}_{3,2}(\frac{1}{2})$ for $i=1,2$
+By Lemma 3.1, if $0\in{\mathscr{W}}_{r}(\alpha)$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$, and there exists $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}(0)$.
 Lemma 3.2.
-Let $\alpha\geq 1$, $q\geq 0$ and $r$ be a positive integer.
-For any $T\in{\mathscr{T}}_{r,\alpha}(q)$, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$, unless $q=0$ and $v$ is the unique source of $T$.
-Proof.
-Since $T\in{\mathscr{T}}_{r,\alpha}(q)$, we have $T\in{\mathscr{T}}_{\alpha}$.
-In the definition of ${\mathscr{T}}_{\alpha}$, the condition at (8) for the mapping $\omega_{\alpha,T}$ from $V(T)$ to $\{0\}\cup{\mathbb{R}}^{+}$ implies that for any vertex $v$ in $T$, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ holds, unless $v$ is the source of $T$ and $q=0$.
-Hence the result follows.
-$\Box$
-4 Proof of Theorem 1.1
-Lemma 4.1.
-Let $\alpha\geq 1$ and $r$ be a positive integer.
-If there exists a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s\leq r$, such that $\sum\limits_{1\leq i\leq s}q_{i}^{-1}=\alpha$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$.
-Proof. By the definition of ${\mathscr{W}}_{r}(\alpha)$ and (8), it is clear that ${\mathscr{T}}_{r,\alpha}(q)\neq\emptyset$ for any $q\in{\mathscr{W}}_{r}(\alpha)$.
-Assume that $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$.
-For each $i\in\llbracket s\rrbracket$, let $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$, $\omega_{i}$ be the mapping $\omega_{\alpha,T_{i}}$ and $z_{i}$ be the unique source of $T_{i}$.
-By definition, $\omega_{i}(z_{i})=q_{i}$ for all $i\in\llbracket s\rrbracket$.
-Now let $T$ be the directed tree obtained from $T_{1},T_{2},\dots,T_{s}$ by adding a new vertex $z$ and adding a new directed edge $(z,z_{i})$ for each $i\in\llbracket s\rrbracket$.
-Let $\omega:V(T)\mapsto{\mathbb{R}}^{+}\cup\{0\}$ be the mapping defined by $\omega(z)=0$ and $\omega(w)=\omega_{i}(w)$ whenever $w\in V(T_{i})$.
-Since $\sum\limits_{1\leq i\leq s}q_{i}^{-1}=\alpha$, condition (8) holds for $\omega$ at $z$.
-For any $v\in V(T)$, if $v\in V(T_{i})$, as $\omega(w)=\omega_{i}(w)$ for all $w\in V(T_{i})$, condition (8) also holds $\omega$ at $v$.
-Since $s\leq r$ and $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket s\rrbracket$, we have $\Delta(T)\leq r$, we have $T\in{\mathscr{T}}_{r,\alpha}(0)$. $\Box$
-Applying the set ${\mathscr{T}}_{r,\alpha}(0)$, we can now obtain the following conclusion for proving Theorem 1.1.
-Lemma 4.2.
-Let $\alpha\geq 1$ and $r$ be a positive integer. For any $T\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
-Proof.
-By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$.
-It suffices to show that $\lambda(T)=\alpha$.
-If $\alpha=1$, then $T\in{\mathscr{T}}_{r,\alpha}(0)$ implies that $|V(T)|=2$, and the result holds.
-Now assume that $\alpha>1$.
-Let $T\in{\mathscr{T}}_{r,\alpha}(0)$.
-By definition, $T$ is a directed tree with a unique source, say $z$.
-By Lemma 3.1, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)\setminus\{z\}$.
-Let $T_{1},T_{2},\dots,T_{s}$ be the components of $T-z$.
-Clearly, $1\leq s\leq r$ and $T_{i}\in{\mathscr{T}}_{\alpha}$ for each $i\in\llbracket s\rrbracket$.
-Let $z_{i}$ be the vertex in $T_{i}$ such that $(z,z_{i})$ is a directed edge in $T$.
-Then, $z_{i}$ the unique source of $T_{i}$ and $q_{i}=\omega_{\alpha,T}(z_{i})$ for each $i\in\llbracket s\rrbracket$.
-By (8), we have
-$0=\omega_{\alpha,T}(z)=\alpha-\sum_{i=1}^{s}q_{i}^{-1}.$ (9)
-In order to prove that $\lambda(T)=\alpha$, by Proposition 2.3, it suffices to show that the mapping $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ defined below satisfies condition (4):
+For any $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
+Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\lambda(T)=\alpha$.
+By definition again, $T$ is a directed tree with a unique source $z$ and $\omega(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$.
+Let $T_{1},T_{2},\dots,T_{s}$ be the components of $T-z$, where $s=od_{T}(z)$.
+Clearly, $1\leq s\leq r$.
+For each $i\in\llbracket s\rrbracket$, $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(\beta_{i})$, where $\omega_{i}$ is the restriction of $\omega$ to $V(T_{i})$, $\beta_{i}=\omega(z_{i})$ and $z_{i}$ is the vertex in $T_{i}$ with $z_{i}\in N^{+}_{T}(z)$.
+Then $z_{i}$ the unique source of $T_{i}$ for each $i\in\llbracket s\rrbracket$ and
+$0=\omega(z)=\alpha-\sum_{i=1}^{s}\omega(z_{i})^{-1}=\alpha-\sum_{i=1}^{s}\beta_{i}^{-1}.$ (7)
+Now let $\phi$ be the mapping from $V(T)$ to ${\mathbb{R}}^{+}$ defined below:
   1. (i).
@@
   2. (ii).
-    for each directed edge $(w_{1},w_{2})\in E(T)$, $\phi(w_{2})=\frac{\phi(w_{1})}{\omega(w_{2})}$, where $\omega$ simply denotes $\omega_{\alpha,T}$.
-We can first show that condition (4) is satisfied for the unique source and all sinks of $T$.
-As $z$ is the only source of $T$ and $N^{+}_{T}(z)=\{z_{i}:i\in\llbracket s\rrbracket\}$, where $N^{+}_{T}(z)$ is the set of out-neighbors of $z$ in $T$, we have
-$\sum_{v\in N_{T}(z)}\phi(v)=\sum_{i=1}^{s}\phi(z_{i})=\sum_{i=1}^{s}\frac{\phi(z)}{\omega(z_{i})}=\phi(z)\sum_{i=1}^{s}q_{i}^{-1}=\alpha\phi(z),$ (10)
+    for each directed edge $(w_{1},w_{2})\in E(T)$, $\phi(w_{2})=\frac{\phi(w_{1})}{\omega(w_{2})}$.
+In order to show that $\lambda(T)=\alpha$, by Proposition 2.2, it suffices to prove the following Claim.
+Claim A: $\phi$ satisfies the condition of (2) for each vertex $u\in V(T)$.
+We first prove Claim A for the case $u=z$ or $u$ is a sink of $T$ (i.e., $N^{+}_{T}(u)=\emptyset$). As $N^{+}_{T}(z)=\{z_{i}:i\in\llbracket s\rrbracket\}$, we have
+$\sum_{v\in N^{+}_{T}(z)}\phi(v)=\sum_{i=1}^{s}\phi(z_{i})=\sum_{i=1}^{s}\frac{\phi(z)}{\omega(z_{i})}=\phi(z)\sum_{i=1}^{s}\beta_{i}^{-1}=\alpha\phi(z),$ (8)
 where the last step follows from (7).
-For any sink $w$ of $T$, $\omega(w)=\alpha$ by (8). If $w^{\prime}$ is the unique in-neighbor of $w$ in $T$, then $\phi(w^{\prime})=\phi(w)\omega(w)$ by the definition of $\phi$, and
-$\sum_{v\in N_{T}(w)}\phi(v)=\phi(w^{\prime})=\phi(w)\omega(w)=\alpha\phi(w).$ (11)
-It remains to consider any vertex $w$ in $T$ which is neither the source nor a sink of $T$.
-Assume that $w^{\prime}$ ie the only in-neighbor of $w$ and $N^{+}_{T}(w)=\{w_{1},w_{2},\dots,w_{t}\}$.
+For the case that $u$ is a sink of $T$, $\omega(u)=\alpha$ by (5). If $u^{\prime}$ is the unique in-neighbor of $u$ in $T$, then $\phi(u^{\prime})=\phi(u)\omega(u)$ by the definition of $\phi$, and
+$\sum_{v\in N_{T}(u)}\phi(v)=\phi(u^{\prime})=\phi(u)\omega(u)=\alpha\phi(u).$ (9)
+Finally, we assume that $u$ is vertex in $T$ which is neither the source nor a sink of $T$.
+Assume that $u^{\prime}$ ie the only in-neighbor of $u$ in $T$ and $N^{+}_{T}(w)=\{u_{1},u_{2},\dots,u_{t}\}$.
 By the definition of $\phi$,
-$\phi(w^{\prime})+\sum_{i=1}^{t}\phi(w_{i})=\omega(w)\phi(w)+\sum_{i=1}^{t}\frac{\phi(w)}{\omega(w_{i})}=\phi(w)\left(\omega(w)+\sum_{i=1}^{t}\omega(w_{i})^{-1}\right)=\alpha\phi(w),$ (12)
-where the last step follows from (8). Thus, $\phi$ satisfies condition (4) for all vertices in $T$, implying that $\lambda(T)=\alpha$ by Proposition 2.3. $\Box$
-Now we are going to establish the following conclusion, by which Theorem 1.1 follows directly.
-Proposition 4.3.
-For any positive numbers $\alpha$ and $r$, where $\alpha\geq 1$ and $r$ is an integer, the following three statements are pairwise equivalent:
-  1. (i).
-    there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$;
-  2. (ii).
-    $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$; and
-  3. (iii).
-    there exist $q_{1},q_{2},\dots,q_{s}\in{\mathscr{W}}_{r}(\alpha)$, where $1\leq s\leq r$, such that $\sum\limits_{i=1}^{s}q_{i}^{-1}=\alpha$.
-Proof. (iii) follows from (ii) directly. By Lemmas 4.2 and 4.1, (i) follows from (iii).
-(i) $\Rightarrow$ (ii) Assume that $T$ is a tree with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
-By Proposition 2.3, there exists a mapping $\phi$ from $V(T)$ to ${\mathbb{R}}^{+}$ such that $\alpha\phi(u)=\sum\limits_{w\in N_{T}(u)}\phi(w)$ for each $u\in V(T)$.
-Then, by Lemma 2.4, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
-Thus, (i) implies (ii).
+$\phi(u^{\prime})+\sum_{i=1}^{t}\phi(u_{i})=\omega(u)\phi(u)+\sum_{i=1}^{t}\frac{\phi(u)}{\omega(u_{i})}=\phi(u)\left(\omega(u)+\sum_{i=1}^{t}\omega(u_{i})^{-1}\right)=\alpha\phi(u),$ (10)
+where the last step follows from (5). Thus, $\phi$ satisfies condition (2) for each $u\in V(T)$, implying that $\lambda(T)=\alpha$ by Proposition 2.2. $\Box$
+By Lemmas 3.1 and 3.2, if $0\in{\mathscr{W}}_{r}(\alpha)$, then there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$, and thus the sufficiency of Theorem 1.1 holds.
+Corollary 3.3.
+If there exist a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r\rrbracket\big{\}}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
+Proof.
+Assume that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, where $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ and $q_{i}>0$ for each $i\in\llbracket r\rrbracket$.
+By Lemma 3.1, there exists $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket r\rrbracket$.
+Assume that $z_{i}$ is the only source of $T_{i}$ for each $i\in\llbracket r\rrbracket$.
+Let $T$ be the directed tree obtained from $T_{1},\cdots,T_{r}$ by adding a new vertex $z$ and directed edges $(z,z_{i})$ for each $i\in\llbracket r\rrbracket$, and let $\omega$ be the mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ defined as follows: $\omega(z)=0$ and $\omega(v)=\omega_{i}(v)$ whenever $v\in V(T_{i})$, where $i\in\llbracket r\rrbracket$.
+By the definition of $T$ and $\omega$, it can be verified easily that $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$.
+Clearly, $r\geq\Delta(T)\geq od_{T}(z)=r$, implying that $\Delta(T)=r$.
+By Lemma 3.2, $\lambda(T)=\alpha$.
 $\Box$
-Clearly, Theorem 1.1 follows from Proposition 4.3 directly. We end this section by a result on the existence of trees $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
-For any directed graph $D$, let $\Delta^{(o)}(D)$ denote the maximum out-degree of $D$. By Lemma 4.2 and its proof, the following conclusion holds.
-Proposition 4.4.
-Let $\alpha\geq 1$ and $r$ be an integer with $\alpha^{2}/4+1<r\leq\alpha^{2}$.
-Assume that there exist a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s\leq r$, such that $\sum\limits_{i=1}^{s}q_{i}^{-1}=\alpha$.
-If $s=r$ or $\Delta^{(o)}(T_{i})=r-1$ for some $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$, where $i\in\llbracket s\rrbracket$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
+Let ${\mathscr{W}}_{r,\alpha}(0)=\big{\{}\omega(v):v\in V(T),(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)\big{\}}$.
 Corollary 3.4.
-For any positive real number $\alpha$ and positive integer $r$ with $\alpha^{2}/4+1<r\leq\alpha^{2}$, if there exist a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket r\rrbracket\}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
+Assume that $0\in{\mathscr{W}}_{r}(\alpha)$.
+If ${\mathscr{W}}_{r,\alpha}(0)$ contains $\beta$ such that $\beta=\alpha-\sum\limits_{i=1}^{r-1}q_{i}^{-1}>0$ holds for a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r-1\rrbracket\big{\}}$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
 4 Proof of Theorem 1.5
-Let $k$ be an integer with $k\geq 2$ and $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$.
-Clearly, for any integer $r$, $r>\frac{1}{4}k+1$ if and only if $r\geq r_{0}$.
-If $k=4m+t$, where $0\leq t\leq 3$, then $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2=m+2$.
-In order to prove Theorem 1.4, by Corollary 4.5, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{r}(\sqrt{k})$) for all integers $s$ in the interval $[r_{0},k]$.
+Let $k\in{\mathbb{N}}^{+}$ with $k\geq 2$ and $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$.
+Clearly, for any $r\in{\mathbb{N}}^{+}$, $r>\frac{1}{4}k+1$ if and only if $r\geq r_{0}$.
+If $k=4m+t$, where $m\in\{0\}\cup{\mathbb{N}}^{+}$ and $0\leq t\leq 3$, then $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2=m+2$.
+In order to prove Theorem 1.5, by Corollary 3.3, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{s}(\sqrt{k})$) for all integers $s$ in the interval $[r_{0},k]$, since this implies that $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}$.
 We start with the following simple result.
@@
   1. (i).
-    When $m=0$ and $1\leq t\leq 3$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[1,t]$; and
+    When $m=0$ and $2\leq t\leq 3$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[2,t]$; and
   2. (ii).
@@
 $\frac{t-1}{\sqrt{t}}=\sqrt{t}-\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t}).$
-Thus, it mains to show that $\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t})$ for $t=3$, i.e., $\frac{1}{\sqrt{3}}\in{\mathscr{W}}_{m+2}(\sqrt{3})$.
-Note that $\frac{2}{\sqrt{3}}\in{\mathscr{W}}_{m+2}(\sqrt{3})$.
-Thus,
-$\frac{\sqrt{3}}{2}=\sqrt{3}-\frac{1}{\frac{2}{\sqrt{3}}}\in{\mathscr{W}}_{m+2}(\sqrt{3}),$
-and
-$\frac{1}{\sqrt{3}}=\sqrt{3}-\frac{1}{\frac{\sqrt{3}}{2}}\in{\mathscr{W}}_{m+2}(\sqrt{3}).$
-(ii) Assume that $m\geq 1$. For any integer $s$ with $3m+t-1\leq s\leq 4m+t$, we have $10\leq 4m+t-s\leq m+$ 1 and
+(ii) Assume that $m\geq 1$. For any $s\in{\mathbb{N}}^{+}$ with $3m+t-1\leq s\leq 4m+t$, we have $10\leq 4m+t-s\leq m+$ 1 and
 $\frac{s}{\sqrt{4m+t}}=\sqrt{4m+t}-\frac{4m+t-s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{k}).$ (11)
 Thus, the result holds. $\Box$
-We are now going to establish the conclusion that for any integer $k=4m+t$, where $m\geq 1$ and $t\in\{0,1,2,3\}$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ holds for all integers $s$ in the interval $[m,4m+t]$, on a case-by-case basis of $t$.
+We are now going to establish the conclusion that for any $k=4m+t$, where $m\in{\mathbb{N}}^{+}$ and $t\in\{0,1,2,3\}$, $\frac{s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+t})$ holds for all integers $s$ in the interval $[m,4m+t]$ by considering each possible value of $t$ separately.
 Proposition 4.2.
-For any positive integer $m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $s$ in $[m,4m]$.
+For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $s$ in $[m,4m]$.
 Proof. We will prove that the conclusion in the following claims. This result is true for integers $s$ in the interval $[3m,4m]$ by Lemma 4.1.
-Claim 1: $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i\geq 1$.
+Claim 1: $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all $i\in{\mathbb{N}}^{+}$.
 By definition, $2\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
 Thus, the claim holds for $i=1$.
-For any positive integer $i$, if $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then by definition,
+For any $i\in{\mathbb{N}}^{+}$, if $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then by definition,
 $\frac{i+2}{i+1}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{i+1}{i}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (12)
@@
 Claim 3: $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i$ with $20\leq i\leq m-$ 2.
-By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. For any integer $i$ with $21\leq i\leq m-$ 2, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
+By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
+For any $i\in{\mathbb{N}}^{+}$ with $21\leq i\leq m-$ 2, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
 $\frac{m-i-1}{m-i}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{m-i}{m-i+1}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (15)
 Thus, the claim holds.
-Claim 4: For any integer $s$ with $m\leq s\leq 4m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
-By Lemma 5.1, Claim 4 holds for $3m-1\leq s\leq 4m$. For any integer $i$ with $20\leq i\leq m-$ 2, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
+Claim 4: For $s\in{\mathbb{N}}^{+}$ with $m\leq s\leq 4m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
+By Lemma 4.1, Claim 4 holds for $3m-1\leq s\leq 4m$. For any $i\in{\mathbb{N}}^{+}$ with $20\leq i\leq m-$ 2, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
 $\frac{2m+i}{2{\sqrt{m}}}=2\sqrt{m}-\frac{m-i-1}{\frac{(m-i-1){\sqrt{m}}}{m-i}}-\frac{i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (16)
 Hence Claim 4 holds for $22m\leq s\leq 3m-$ 2. In particular, $\sqrt{m}=\frac{2m}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
-For any integer $i$ with $10\leq i\leq m-$ 1, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
+For any $i\in{\mathbb{N}}^{+}$ with $10\leq i\leq m-$ 1, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
 $\frac{2{\sqrt{m}}\,(m-i-1)}{2m-i}=2\sqrt{m}-\frac{m+1}{\frac{(2m-i)}{2\sqrt{m}}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ (17)
@@
 Proposition 4.3.
-For any positive integer $m$, $\frac{s}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for all integers $s$ in $[m,4m+1]$.
+For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for all integers $s$ in $[m,4m+1]$.
 Proof. We first prove the following claims.
-Claim 1: For all $i\geq 1$, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ and $\frac{(2i+1)m}{i\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
+Claim 1: For any $i\in{\mathbb{N}}^{+}$, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ and $\frac{(2i+1)m}{i\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
 Let $Q(i)=\frac{(2i+1)m}{i\sqrt{4m+1}}$. By Lemma 4.1, $Q(1)=\frac{3m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ By definition, if $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then,
@@
 Thus, Claim 1 follows.
-Claim 2: For any integer $i$ with $1\leq i\leq m$, we have $\frac{3m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
+Claim 2: For any $i\in\llbracket m\rrbracket$, we have $\frac{3m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
 By Claim 1, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for $i\geq 1$. If $1\leq i\leq m$, then
@@
 Claim 2 follows.
-Claim 3: For any integer $i$ with $0\leq i\leq m$, we have $\frac{2m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
+Claim 3: For any $i\in\{0\}\cup\llbracket m\rrbracket$, we have $\frac{2m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
 Let $P(i)=\frac{2m-i}{\sqrt{4m+1}}$. By Claim 2, $P(0)=\frac{2m}{\sqrt{4m+1}}=\frac{3m-m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. By definition,
@@
 Hence the claim holds for $i=0,1$.
-It suffices to show that for any integer $i$ with $20\leq i\leq m-$ 2, if $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
+It suffices to show that for any $i\in\{0\}\cup\llbracket m-2\rrbracket$, if $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
 Assume that $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then
@@
 Proposition 4.4.
-For any positive integer $m$, $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for all integers $s$ in $[m,4m+2]$.
+For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for all integers $s$ in $[m,4m+2]$.
 Proof. We first prove the following claims.
-Claim 1: For any integer $i\geq 1$, both $\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$ belong to ${\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
+Claim 1: For any $i\in{\mathbb{N}}^{+}$, both $\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$ belong to ${\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
 Let $P(i)=\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $Q(i)=\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$.
@@
 Thus, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for each integer $i$ with $-2\leq i\leq 1$.
-It suffices to show that for any $i$ with $10\leq i\leq m-$ 1, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ implies that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
-Observe that for any $i$ with $10\leq i\leq m-$ 1,
+It suffices to show that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ implies that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
+Observe that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$,
 $b:=\sqrt{4m+2}-\frac{m+1}{W(i)}=\sqrt{4m+2}-\frac{m+1}{\frac{2m-i}{\sqrt{4m+2}}}=\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (33)
@@
 By Claim 2, $\frac{2m+i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for $i=0,1$.
-For any integer $i$ with $20\leq i\leq m-$ 2, it is shown in (35) that $\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ Thus,
+For any $i\in\{0\}\cup\llbracket m-2\rrbracket$, it is shown in (33) that $\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ Thus,
 $\frac{2m+2+i}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m-i-1}{\frac{(m-i-1)\cdot\sqrt{4m+2}}{2m-i}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (35)
@@
 Proposition 4.5.
-For any positive integer $m$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $s$ in $[m,4m+3]$.
+For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $s$ in $[m,4m+3]$.
 Proof. We first prove the following claims.
-Claim 1: For any positive integer $i$, $\frac{(i+1)\sqrt{4m+3}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ and $\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
+Claim 1: For any $i\in{\mathbb{N}}^{+}$, $\frac{(i+1)\sqrt{4m+3}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ and $\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
 Let $f_{1}(i)=\frac{(i+1)\sqrt{4m+3}}{2i+1}$ and $f_{2}(i)=\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}$. Obviously, $\sqrt{4m+3}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$,
@@
 implying that $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each $i\in\{-2,-1,0\}$.
-It remains to show that for any integer $i$ with $10\leq i\leq m-$ 1, $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ implies that $g(i+1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Observe that
+It remains to show that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$, $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ implies that $g(i+1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Observe that
 $\sqrt{4m+3}-\frac{m+1}{g(i)}=\sqrt{4m+3}-\frac{(m+1)}{\frac{2m-i}{\sqrt{4m+3}}}=\frac{(m-i-1)\sqrt{4m+3}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (47)
@@
 Hence Claim 2 follows.
-Claim 3: For any integer $s$ with $32m\leq s\leq 4m+$ 3, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
+Claim 3: For any $s\in{\mathbb{N}}^{+}$ with $32m\leq s\leq 4m+$ 3, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
 By Lemma 4.1, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $33m+2\leq s\leq 4m+$ 3.
@@
 By Claim 2, $h(0),h(1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
-For any integer $i$ with $0\leq i\leq m$, by Claim 1, $\frac{(m-i+1)\sqrt{4m+3}}{2(m-i)+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$, implying that
+For any $i\in\{0\}\cup\llbracket m\rrbracket$, by Claim 1, $\frac{(m-i+1)\sqrt{4m+3}}{2(m-i)+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$, implying that
 $h(i+2)=\frac{2m+2+i}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{(m+1-i)}{\frac{(m-i+1)\sqrt{4m+3}}{2\cdot(m-i)+1}}-\frac{i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (49)
@@
 Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$.
-By Lemma 5.1 and Propositions 5.25.35.4 and 5.5, for each integer $s$ with $\left\lfloor{\frac{1}{4}k}\right\rfloor\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ holds, implying that
+By Lemma 4.1 and Propositions 4.24.34.4 and 4.5, for each integer $s$ with $r_{0}\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})\subseteq{\mathscr{W}}_{s}(\sqrt{k})$ holds.
+Note that
 $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}.$ (50)
-Then, by Corollary 4.5, for any integer $r$ in $[r_{0},k]$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$. $\Box$
-6 Problems and Remarks
+By Corollary 3.3, for any integer $s$ in $[r_{0},k]$, there exists a tree $T$ with $\Delta(T)=s$ and $\lambda(T)=\sqrt{k}$. $\Box$
+5 Further study
 Recall that $\Lambda$ is the set of $\lambda(T)$ over all trees $T$ of order at least $2$.
@@
 Now let $\Lambda^{\prime}$ be the set of real numbers $\alpha$ such that $\alpha$ is the largest real root of some $P\in{\cal P}$.
-Problem 6.1.
+Problem 5.1.
 Is there a positive real number $c$ such that $\{\alpha\in\Lambda^{\prime}:\alpha\geq c\}\subseteq\Lambda$?
-It can be proved by applying Proposition 4.4 that, for any positive integers $p$ and $q$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if $4|(p-q)$, $\sqrt{p+\sqrt{q}}\in\Lambda$ if $p\geq q$, and $\frac{\sqrt{2(2p+1)+2\sqrt{4q+1}}}{2}\in\Lambda$ if $q\leq p^{2}+p$.
+It can be proved by applying Theorem 1.1 that, for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if $4|(p-q)$, $\sqrt{p+\sqrt{q}}\in\Lambda$ if $p\geq q$, and $\frac{\sqrt{(2p+1)+\sqrt{4q+1}}}{\sqrt{2}}\in\Lambda$ if $q\leq p^{2}+p$.
 As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ whenever $4|(p-q)$.
 Proposition 5.2.
-For any positive integers $p$ and $q$ with $p\geq q$, if $p-q$ is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\{q,\frac{p-q}{4}\}$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
+For any $p,q\in{\mathbb{N}}^{+}$ with $p\geq q$, if $p-q$ is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\left\{q,\frac{p-q}{4}\right\}$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
 Proof. Assume that $p-q=4d$. Let $r=\max\{q,d\}+1$ and $\alpha=\frac{\sqrt{p}+\sqrt{q}}{2}$. By definition,
@@
 $d\cdot\left(\frac{\sqrt{p}-\sqrt{q}}{2}\right)^{-1}=\frac{\frac{p-q}{4}}{\frac{\sqrt{p}-\sqrt{q}}{2}}=\frac{\sqrt{p}+\sqrt{q}}{2}=\alpha.$ (53)
-By Proposition 4.4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
+By Corollary 3.4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
 We wonder if $4|(p-q)$ is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\Lambda$.
-For any positive integers $s$ and $t$, $\frac{s+t}{2}=\frac{\sqrt{s^{2}}+\sqrt{t^{2}}}{2}$.
-As $P(G,\lambda)\in{\cal P}$, each rational root of $P(G,\lambda)=0$ is an integer.
-It follows that $\frac{s+t}{2}\in\Lambda$ if and only if $s$ and $t$ have the same parity, i.e., $4|(s^{2}-t^{2})$.
-Problem 6.3.
-Is it true that for any positive integers $p$ and $q$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if and only if $p-q$ is a multiple of $4$?
+Problem 5.3.
+Is it true that for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if and only if $4\big{|}(p-q)$?
 Let $T$ be any tree with $\alpha=\lambda(T)$ and $r=\Delta(T)$.
@@
 By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$.
-By Theorem 1.1, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
+From the proving of Lemma 2.3, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
 We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
-Problem 6.4.
-Let $\lambda\in\Lambda$ and $r$ be an integer with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?
+Problem 5.4.
+Let $\lambda\in\Lambda$ and $r\in{\mathbb{N}}^{+}$ with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?
+Note that some similar work on Laplacian spectral radii of trees has been completed in [8].
+Conflict of interest: The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Data availability statements: Not applicable.
 Acknowledgement

### v3

@@
-An algebraic characterization for spectral radii of trees
+Existence of trees with prescribed maximum degrees and spectral radii
 Abstract
-There have been various results on spectral radii of trees, studying the relations between spectral radius $\lambda(T)$ of a tree $T$ and some other parameters of $T$.
-In this article, we provide an algebraic characterization for the spectral radii of trees, and apply this new result to study the existence of trees with given spectral radius and maximum degree.
-For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set defined recursively: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,1\leq i\leq s\}$, if $\beta:=\alpha-\sum\limits_{i=1}^{s}q_{i}^{-1}\geq 0$ and $s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$.
-We first show that for any positive integer $r$ and real number $\alpha\geq 1$, there exists a tree $T$ with maximum degree $\Delta(T)\leq r$ and spectral radius $\lambda(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
-It follows that the set of spectral radii of non-trivial trees is exactly the set of positive numbers $\alpha$ such that $0\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$.
-Applying this conclusion, we then prove that for any positive integers $r\geq 2$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
+It is well known that the spectral radius $\rho(T)$ of a tree $T$ with at least $3$ vertices has the property that $\frac{1}{4}\rho(T)^{2}+1<\Delta(T)\leq\rho(T)^{2}$, where $\Delta(T)$ is the maximum degree of $T$.
+Let $\mathbb{P}$ denote the set of spectral radii of all non-trivial trees.
+In this article, we study the inverse problem that for any $\alpha\in\mathbb{P}$ and integer $r$ satisfying the condition $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, is there a tree $T$ such that $\Delta(T)=r$ and $\rho(T)=\alpha$?
+For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote a set of non-negative real numbers defined as follows: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,1\leq i\leq s\}$, if $\beta:=\alpha-\sum\limits_{i=1}^{s}q_{i}^{-1}\geq 0$ and $s\leq r-\left\lceil\frac{\beta}{\beta+1}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$.
+We first show that $0\in{\mathscr{W}}_{r}(\alpha)$ if and only if there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$.
+It follows directly that $\mathbb{P}$ is exactly the set of positive numbers $\alpha$ such that $0\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$.
+Applying this conclusion, we prove that for any two positive integers $r\geq 2$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
 AMS Subject classification: 05C05; 05C20; 05C50
@@
 1 Introduction
-All graphs considered in this article are simple and undirected.
-For a graph $G$, let $V(G)$, $E(G)$ and $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively.
-For any $u\in V(G)$, let $N_{G}(u)$ (or simply $N(u)$) denote the set of neighbours of $u$ in $G$, and let $d_{G}(u)$ or (simply $d(u)$) be the size of $N_{G}(u)$, called the degree of $u$ in $G$.
-Let $L(G)$ denote the set of vertices $v\in V(G)$ with $d_{G}(v)=1$.
-Let ${\mathbb{N}}^{+}$ denote the set of positive integers and for any $k\in{\mathbb{N}}^{+}$, let $\llbracket k\rrbracket=\{i\in{\mathbb{N}}^{+}:i\leq k\}$.
-For a graph $G=(V,E)$, where $V=\{v_{i}:i\in\llbracket k\rrbracket\}$, its adjacency matrix $A(G)$ according to the vertex ordering $v_{1},v_{2},\ldots,v_{n}$ is defined to be the $0-1$ matrix $(a_{i,j})_{n\times n}$, where $a_{i,j}=1$ if and only if $v_{i}v_{j}\in E$.
-The characteristic polynomial of $G$, denoted by $P(G,\lambda)$, is defined to be the function $det(\lambda I_{n}-A(G))$, where $I_{n}$ is the identical matrix of size $n$.
-It is known that $P(G,\lambda)$ is independent of the vertex ordering $v_{1},v_{2},\dots,v_{n}$, and the roots of the equation $P(G,\lambda)=0$, called the eigenvalues of $G$, are real numbers (see [6]), and the largest eigenvalue of $G$, denoted by $\lambda(G)$, is called the spectral radius of $G$.
+All graphs in this article are finite, simple and undirected.
+For a graph $G$, let $V(G)$, $E(G)$ $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively.
+If $|V(G)|=n$, then $G$ has exactly $n$ eigenvalues, and they are all real numbers (see [6]).
+The largest eigenvalue of $G$, denoted by $\rho(G)$, is called the spectral radius of $G$.
 The spectral radii of graphs, which have been studied extensively (see a survey [33] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 11, 18, 19, 29, 30, 33]).
-The distribution of spectral radii among different graphs is a topic of ongoing research with connections to various graph properties and invariants (see [7, 12, 13, 16, 20, 22, 23]).
+The distribution of spectral radii among different graphs is a topic of ongoing research with connections to various graph properties and invariants (see [7, 9, 16, 18, 19, 23, 27, 29, 30]).
 In this article, we focus on the study of spectral radii of trees.
-There are various known results on the relation between $\lambda(T)$ and other parameters of a tree $T$.
-Guo and Tan [11] found an upper bound of $\lambda(T)$ in terms of the order and matching number of $T$.
-Lovász and Pelikán [15] determined the maximal and minimal value of $\lambda(T)$ among all trees $T$ of order $n$, Simić and Tos̆ić [24] found the tree that has the largest spectral radius among all trees with given order and maximum degree, and Stevanović [25] provided a relation between $\lambda(T)$ and $\Delta(T)$: for any tree $T$ with $|V(T)|\geq 3$,
-$\frac{1}{4}\lambda(T)^{2}+1<\Delta(T)\leq\lambda(T)^{2}.$ (1)
-In this article, we will first provide a characterization on $r$ and $\alpha$ for the existence of trees $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
-Let $\Lambda$ denote the set of $\lambda(T)$ for all non-trivial trees $T$ (i.e., trees of order at least $2$).
-Clearly, $\alpha\geq 1$ for each $\alpha\in\Lambda$.
-Let ${\mathbb{R}}^{+}$ denote the set of positive real numbers, and for any $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of non-negative real numbers defined below:
+There are various known results on the relation between $\rho(T)$ and other parameters of a tree $T$.
+Lovász and Pelikán [22] determined the maximal and minimal value of $\rho(T)$ over all trees $T$ with a fixed order, Guo and Tan [13] found an upper bound of $\rho(T)$ in terms of the order and matching number of $T$, and Simić and Tos̆ić [31] confirmed the tree that has the largest spectral radius over all trees with given order and maximum degree.
+Stevanović [32] provided a relation between $\rho(T)$ and $\Delta(T)$, as stated below.
+Theorem 1 ([32]).
+For any tree $T$ with $|V(T)|\geq 3$, $\frac{1}{4}\rho(T)^{2}+1<\Delta(T)\leq\rho(T)^{2}.$
+A graph $G$ is said to be non-trivial if $|V(G)|\geq 2$. Let $\mathbb{P}$ denote the set of $\rho(T)$ over all non-trivial trees $T$. In this article, we will study the following inverse problem:
+Problem 1.
+Is it true that for any $\alpha\in\mathbb{P}$ and any integer $r$ satisfying the condition $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, there exists a tree $T$ with $\rho(T)=\alpha$ and $\Delta(T)=r$?
+The above inverse problem connects to characterizing possible spectra of graphs and is similar in spirit to the inverse eigenvalue problem for general matrices.
+Studying inverse problems in graph theory is important because it offers a reverse perspective on fundamental questions, connects structure to invariants, underpins classical existence theorems, inspires constructive algorithms, etc.
+In fact, many famous problems in graph theory are inverse in nature.
+For example, the reconstruction conjecture in graph theory (see [34]) says that any connected graph with at least three vertices is determined uniquely by its subgraphs obtained by removing one vertex, the realization problem of degree sequences studies sequences of positive integers which are degree sequences of graphs (see Erdős and Gallai [10], Havel [15] and Hakimi [14]), the inverse spectral problem studies possible sets of real numbers which are spectra of graphs (see [17]), and the inverse Wiener problem studies whether a given integer can be the Wiener index of some graph (see [20, 35, 36]).
+Let $\mathbb{N}$$^{+}$ and $\mathbb{R}$$^{+}$ denote the set of positive integers and the set of positive real numbers, respectively.
+For any $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of non-negative real numbers defined below:
   1. (i).
@@
 It can be verified that ${\mathscr{W}}_{1}(1)=\{0,1\}$, ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha\in{\mathbb{R}}^{+}$ with $\alpha\neq 1$, ${\mathscr{W}}_{2}(\sqrt{2})=\left\{0,\sqrt{2},(\sqrt{2})^{-1}\right\}$ and ${\mathscr{W}}_{2}(2)=\big{\{}\frac{i+1}{i}:i\geq 1\big{\}}$.
-Clearly, the property ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ holds for any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$.
-Our first aim in this article is to establish a connection between the existence of a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ and the existence of 00 in the set ${\mathscr{W}}_{r}(\alpha)$.
-Theorem 1.1.
-For any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$, there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
-A characterization of the set $\Lambda$ follows from Theorem 1.1 and (1) directly.
-Corollary 1.2.
-$\Lambda=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)\big{\}}$.
+Clearly, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ holds for any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$.
+Our first aim in this article is to establish the following result, providing a connection between the existence of a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$ and whether $0\in{\mathscr{W}}_{r}(\alpha)$.
+Theorem 2.
+For any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$, there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
+A characterization of the set $\mathbb{P}$ follows directly from Theorems 1 and 2.
+Corollary 1.
+$\mathbb{P}=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)\big{\}}$.
 If $r$ is removed from the definition of the set ${\mathscr{W}}_{r}(\alpha)$, we can define the set ${\mathscr{W}}(\alpha)$ as follows:
@@
 Clearly, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}(\alpha)$ for any $r\in{\mathbb{N}}^{+}$, and the following conclusion holds.
-Corollary 1.3.
-$\Lambda=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}(\alpha)\big{\}}$.
-The conclusion of (1) tells that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$ is a necessary condition for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
-We wonder if this condition is also sufficient.
-Conjecture 1.4.
-Let $\alpha\in\Lambda\setminus\{1\}$. If $r$ is an integer such that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, then there exists a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
-Our second aim is to apply Theorem 1.1 to prove Conjecture 1.4 for the case $\alpha=\sqrt{k}$ for each $k\in{\mathbb{N}}^{+}\setminus\{1\}$.
-Certainly, it includes the case $\alpha=k$ for each $k\in{\mathbb{N}}^{+}\setminus\{1\}$.
-Theorem 1.5.
-For any $k,r\in{\mathbb{N}}^{+}\setminus\{1\}$, there exists a tree $T$ with $\lambda(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
+Corollary 2.
+$\mathbb{P}=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}(\alpha)\big{\}}$.
+Our second aim in this article is to apply Theorem 2 to prove that Problem 1 has a positive answer for the case $\alpha=\sqrt{k}$ for any $k\in{\mathbb{N}}^{+}\setminus\{1\}$.
+Certainly, it includes the case $\alpha=k\in{\mathbb{N}}^{+}\setminus\{1\}$.
+Theorem 3.
+For any $k,r\in{\mathbb{N}}^{+}\setminus\{1\}$, there exists a tree $T$ with $\rho(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
 The remainder of this article is structured as follows.
-In Sections 2 and 3, we prove Theorem 1.1’s necessity and sufficiency, respectively.
-In Section 4, we complete the proof of Theorem 1.5.
+In Sections 2 and 3, we prove Theorem 2's necessity and sufficiency, respectively.
+In Section 4, we complete the proof of Theorem 3.
 Finally, in Section 5, we provide some open problems arising from our work.
-2 The necessity of Theorem 1.1
+2 The necessity of Theorem 2
 The Perron–Frobenius theorem, proved by Oskar Perron [26] in 1907 and Georg Frobenius [12] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices.
-Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectral radius $\lambda(G)$ of a connected graph $G$ is obtained.
+Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectral radius $\rho(G)$ of a connected graph $G$ is obtained.
 Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$.
-Then, $\lambda(G)=\lambda_{1}(G)$.
-We say $G$ is non-trivial if $|V(G)|\geq 2$.
-Theorem 2.1 (Perron–Frobenius theorem).
+Then, $\rho(G)=\lambda_{1}(G)$.
+Theorem 4 (Perron–Frobenius theorem).
 Let $G$ be a connected non-trivial graph.
@@
 Furthermore, $\lambda_{1}(G)$ is the only eigenvalue of $G$ with this property.
-Applying Theorem 2.1, the next conclusion follows directly.
-Proposition 2.2.
-Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number.
-Then, $\lambda(G)=\alpha$ if and only if there is a mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
+Applying Theorem 4, the next conclusion follows directly.
+Proposition 1.
+Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number. Then, $\rho(G)=\alpha$ if and only if there is a mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
 $\alpha\phi(u)=\sum_{w\in N(u)}\phi(w).$ (1)
-Note that for any mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$, the condition of (2) holds if and only if $\phi$ is an eigenvector of $A(G)$ with respect to $\lambda(G)$.
-Now we are going to prove the following lemma by which the necessity of Theorem 1.1 follows directly.
-Lemma 2.3.
-Let $T$ be a non-trivial tree and $\alpha=\lambda(T)$, and let $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ be an eigenvector of $A(T)$ with respect to $\alpha$.
+Note that for any mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$, the condition of (1) holds if and only if $\phi$ is an eigenvector of $A(G)$ with respect to $\rho(G)$.
+Now we are going to prove the following lemma by which the necessity of Theorem 2 follows directly.
+Lemma 1.
+Let $T$ be a non-trivial tree and $\alpha=\rho(T)$, and let $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ be an eigenvector of $A(T)$ with respect to $\alpha$.
 Then both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$, where $r=\Delta(T)$.
-Proof. Since $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is an eigenvector of $A(T)$ with respect to $\alpha=\lambda(T)$, (2) holds for each $u\in V(T)$. Let’s first prove the following claim.
+Proof. Since $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is an eigenvector of $A(T)$ with respect to $\alpha=\rho(T)$, (1) holds for each $u\in V(T)$. Let's first prove the following claim.
 Claim 1: For an arbitrary vertex $z$ in $T$, if $T^{\prime}$ is the orientation of $T$ with $z$ as its source, as shown in Figure 1, then $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$ holds for every directed edge $(w_{1},w_{2})$ in $T^{\prime}$.
@@
 $\Box$
-Proof of the necessity of Theorem 1.1.
-We follow the conditions in Lemma 2.3.
+Proof of the necessity of Theorem 2.
+We follow the conditions in Lemma 1.
 Take any edge $u_{1}u_{2}\in E(T)$, where $u_{1}$ is a leaf of $T$.
-We have $\alpha\phi(u_{1})=\phi(u_{2})$ by (2), implying that $\alpha^{-1}=\frac{\phi(u_{2})}{\phi(u_{1})}\in{\mathscr{W}}_{r}(\alpha)$ by Lemma 2.3.
+We have $\alpha\phi(u_{1})=\phi(u_{2})$ by (1), implying that $\alpha^{-1}=\frac{\phi(u_{2})}{\phi(u_{1})}\in{\mathscr{W}}_{r}(\alpha)$ by Lemma 1.
 Thus, $0=\alpha-\frac{1}{\alpha^{-1}}\in{\mathscr{W}}_{r}(\alpha)$ by definition.
-Hence the result holds. $\Box$
-3 The sufficiency of Theorem 1.1
+$\Box$
+3 The sufficiency of Theorem 2
 In this section, we always assume that $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$.
@@
 By the definitions of ${\mathscr{W}}_{r}(\alpha)$ and ${\mathscr{T}}_{r,\alpha}(\beta)$, the following conclusion holds.
-Lemma 3.1.
+Lemma 2.
 For any $\beta\in{\mathscr{W}}_{r}(\alpha)$, ${\mathscr{T}}_{r,\alpha}(\beta)\neq\emptyset$.
@@
 $\frac{1}{2}$1222
-[…]
+$\frac{1}{2}$$\frac{2}{3}$$\frac{3}{2}$$\frac{3}{2}22$
     (a) $(T_{1},\omega_{1})$                 (b) $(T_{2},\omega_{2})$
 Figure 2: $(T_{i},\omega_{i})\in\mathcal{J}_{3,2}(\frac{1}{2})$ for $i=1,2$
-By Lemma 3.1, if $0\in{\mathscr{W}}_{r}(\alpha)$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$, and there exists $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}(0)$.
-Lemma 3.2.
-For any $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
-Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\lambda(T)=\alpha$.
+By Lemma 2, if $0\in{\mathscr{W}}_{r}(\alpha)$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$, and there exists $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}(0)$.
+Lemma 3.
+For any $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\rho(T)=\alpha$.
+Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\rho(T)=\alpha$.
 By definition again, $T$ is a directed tree with a unique source $z$ and $\omega(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$.
-Let $T_{1},T_{2},\dots,T_{s}$ be the components of $T-z$, where $s=od_{T}(z)$.
+Let $T_{1},T_{2},\dots,T_{s}$ be the components of TzT-z, where $s=od_{T}(z)$.
 Clearly, $1\leq s\leq r$.
@@
     for each directed edge $(w_{1},w_{2})\in E(T)$, $\phi(w_{2})=\frac{\phi(w_{1})}{\omega(w_{2})}$.
-In order to show that $\lambda(T)=\alpha$, by Proposition 2.2, it suffices to prove the following Claim.
+In order to show that $\rho(T)=\alpha$, by Proposition 1, it suffices to prove the following Claim.
 Claim A: $\phi$ satisfies the condition of (1) for each vertex $u\in V(T)$.
@@
 $\phi(u^{\prime})+\sum_{i=1}^{t}\phi(u_{i})=\omega(u)\phi(u)+\sum_{i=1}^{t}\frac{\phi(u)}{\omega(u_{i})}=\phi(u)\left(\omega(u)+\sum_{i=1}^{t}\omega(u_{i})^{-1}\right)=\alpha\phi(u),$ (9)
-where the last step follows from (5). Thus, $\phi$ satisfies condition (2) for each $u\in V(T)$, implying that $\lambda(T)=\alpha$ by Proposition 2.2. $\Box$
-By Lemmas 3.1 and 3.2, if $0\in{\mathscr{W}}_{r}(\alpha)$, then there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$, and thus the sufficiency of Theorem 1.1 holds.
-Corollary 3.3.
-If there exist a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r\rrbracket\big{\}}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
+where the last step follows from (4). Thus, $\phi$ satisfies condition (1) for each $u\in V(T)$, implying that $\rho(T)=\alpha$ by Proposition 1. $\Box$
+By Lemmas 2 and 3, if $0\in{\mathscr{W}}_{r}(\alpha)$, then there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$, and thus the sufficiency of Theorem 2 holds.
+Corollary 3.
+If there exist a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r\rrbracket\big{\}}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\alpha$.
 Proof.
 Assume that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, where $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ and $q_{i}>0$ for each $i\in\llbracket r\rrbracket$.
-By Lemma 3.1, there exists $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket r\rrbracket$.
+By Lemma 2, there exists $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket r\rrbracket$.
 Assume that $z_{i}$ is the only source of $T_{i}$ for each $i\in\llbracket r\rrbracket$.
@@
 Clearly, $r\geq\Delta(T)\geq od_{T}(z)=r$, implying that $\Delta(T)=r$.
-By Lemma 3.2, $\lambda(T)=\alpha$.
+By Lemma 3, $\rho(T)=\alpha$.
 $\Box$
 Let ${\mathscr{W}}_{r,\alpha}(0)=\big{\{}\omega(v):v\in V(T),(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)\big{\}}$.
-Corollary 3.4.
+Corollary 4.
 Assume that $0\in{\mathscr{W}}_{r}(\alpha)$.
-If ${\mathscr{W}}_{r,\alpha}(0)$ contains $\beta$ such that $\beta=\alpha-\sum\limits_{i=1}^{r-1}q_{i}^{-1}>0$ holds for a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r-1\rrbracket\big{\}}$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
-4 Proof of Theorem 1.5
+If ${\mathscr{W}}_{r,\alpha}(0)$ contains $\beta$ such that $\beta=\alpha-\sum\limits_{i=1}^{r-1}q_{i}^{-1}>0$ holds for a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r-1\rrbracket\big{\}}$, then there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\alpha$.
+4 Proof of Theorem 3
 Let $k\in{\mathbb{N}}^{+}$ with $k\geq 2$ and $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$.
@@
 If $k=4m+t$, where $m\in\{0\}\cup{\mathbb{N}}^{+}$ and $0\leq t\leq 3$, then $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2=m+2$.
-In order to prove Theorem 1.5, by Corollary 3.3, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{s}(\sqrt{k})$) for all integers $s$ in the interval $[r_{0},k]$, since this implies that $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}$.
+In order to prove Theorem 3, by Corollary 3, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{s}(\sqrt{k})$) for all integers $s$ in the interval $r_{0},k]$, since this implies that $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}$.
 We start with the following simple result.
-Lemma 4.1.
+Lemma 4.
 Let $k=4m+t\geq 2$, where $m$ and $t$ are integers with $m\geq 0$ and $0\leq t\leq 3$.
@@
   2. (ii).
-    when $m\geq 1$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[3m+t-1,4m+t]$.
+    when $m\geq 1$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval [3m+t1,4m+t][3m+t-1,4m+t].
 Proof. (i). As $m=0$, we have $k=t$ and thus, $\sqrt{k}=\sqrt{t}$. By definition, $\frac{t}{\sqrt{t}}=\sqrt{t}\in{\mathscr{W}}_{m+2}(\sqrt{t})$ and for $2\leq t\leq 3$,
 $\frac{t-1}{\sqrt{t}}=\sqrt{t}-\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t}).$
-(ii) Assume that $m\geq 1$. For any $s\in{\mathbb{N}}^{+}$ with $3m+t-1\leq s\leq 4m+t$, we have $10\leq 4m+t-s\leq m+$ 1 and
+(ii) Assume that $m\geq 1$. For any $s\in{\mathbb{N}}^{+}$ with $3m+t-1\leq s\leq 4m+t$, we have $0\leq 4m+t-s\leq m+1$ and
 $\frac{s}{\sqrt{4m+t}}=\sqrt{4m+t}-\frac{4m+t-s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{k}).$ (10)
@@
 We are now going to establish the conclusion that for any $k=4m+t$, where $m\in{\mathbb{N}}^{+}$ and $t\in\{0,1,2,3\}$, $\frac{s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+t})$ holds for all integers $s$ in the interval $[m,4m+t]$ by considering each possible value of $t$ separately.
-Proposition 4.2.
+Proposition 2.
 For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $s$ in $[m,4m]$.
-Proof. We will prove that the conclusion in the following claims. This result is true for integers $s$ in the interval $[3m,4m]$ by Lemma 4.1.
+Proof. We will prove that the conclusion in the following claims. This result is true for integers $s$ in the interval $[3m,4m]$ by Lemma 4.
 Claim 1: $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all $i\in{\mathbb{N}}^{+}$.
@@
 Thus, Claim 2 holds.
-Claim 3: $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i$ with $20\leq i\leq m-$ 2.
+Claim 3: $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i$ with $0\leq i\leq m-2$.
 By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
-For any $i\in{\mathbb{N}}^{+}$ with $21\leq i\leq m-$ 2, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
+For any $i\in{\mathbb{N}}^{+}$ with $1\leq i\leq m-2$, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
 $\frac{m-i-1}{m-i}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{m-i}{m-i+1}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (14)
@@
 Claim 4: For $s\in{\mathbb{N}}^{+}$ with $m\leq s\leq 4m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
-By Lemma 4.1, Claim 4 holds for $3m-1\leq s\leq 4m$. For any $i\in{\mathbb{N}}^{+}$ with $20\leq i\leq m-$ 2, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
+By Lemma 4, Claim 4 holds for $3m-1\leq s\leq 4m$. For any $i\in{\mathbb{N}}^{+}$ with $0\leq i\leq m-2$, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
 $\frac{2m+i}{2{\sqrt{m}}}=2\sqrt{m}-\frac{m-i-1}{\frac{(m-i-1){\sqrt{m}}}{m-i}}-\frac{i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (15)
-Hence Claim 4 holds for $22m\leq s\leq 3m-$ 2. In particular, $\sqrt{m}=\frac{2m}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
-For any $i\in{\mathbb{N}}^{+}$ with $10\leq i\leq m-$ 1, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
+Hence Claim 4 holds for $2m\leq s\leq 3m-2$. In particular, $\sqrt{m}=\frac{2m}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
+For any $i\in{\mathbb{N}}^{+}$ with $0\leq i\leq m-1$, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
 $\frac{2{\sqrt{m}}\,(m-i-1)}{2m-i}=2\sqrt{m}-\frac{m+1}{\frac{(2m-i)}{2\sqrt{m}}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ (16)
@@
 By Claim 4, the conclusion holds. $\Box$
-Proposition 4.3.
+Proposition 3.
 For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for all integers $s$ in $[m,4m+1]$.
@@
 Claim 1: For any $i\in{\mathbb{N}}^{+}$, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ and $\frac{(2i+1)m}{i\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
-Let $Q(i)=\frac{(2i+1)m}{i\sqrt{4m+1}}$. By Lemma 4.1, $Q(1)=\frac{3m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ By definition, if $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then,
+Let $Q(i)=\frac{(2i+1)m}{i\sqrt{4m+1}}$. By Lemma 4, $Q(1)=\frac{3m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ By definition, if $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then,
 $b:=\sqrt{4m+1}-\frac{m}{Q(i)}=\sqrt{4m+1}-\frac{m}{\frac{(2i+1)m}{i\sqrt{4m+1}}}=\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ (18)
@@
 implying that $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ by definition. Thus, Claim 3 holds.
-By Lemma 4.1 and Claims 2 and 3, the conclusion holds. $\Box$
-Proposition 4.4.
+By Lemma 4 and Claims 2 and 3, the conclusion holds. $\Box$
+Proposition 4.
 For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for all integers $s$ in $[m,4m+2]$.
@@
 Let $P(i)=\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $Q(i)=\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$.
-By Lemma 4.1, $P(1)=\frac{3m+1}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ and $Q(1)=\frac{3m+2}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
+By Lemma 4, $P(1)=\frac{3m+1}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ and $Q(1)=\frac{3m+2}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
 Thus, Claim 1 holds for $i=1$.
@@
 Thus, Claim 1 holds.
-Claim 2: $\frac{2m-i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $i$ with $-2\leq i\leq m+1$.
+Claim 2: $\frac{2m-i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $i$ with $2\leq i\leq m+1$.
 Let $W(i)=\frac{2m-i}{\sqrt{4m+2}}$. By Claim 1, $W(-2)=\frac{(2m+1)m+m}{m\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. By definition,
@@
 $W(1)=\frac{2m-1}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{1}{\frac{2m+1}{\sqrt{4m+2}}}-\frac{m}{\frac{2m}{\sqrt{4m+2}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (31)
-Thus, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for each integer $i$ with $-2\leq i\leq 1$.
+Thus, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for each integer $i$ with $2\leq i\leq 1$.
 It suffices to show that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ implies that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
@@
 Hence Claim 3 holds.
-By Lemma 4.1 and Claims 2 and 3, the conclusion holds. $\Box$
-Proposition 4.5.
+By Lemma 4 and Claims 2 and 3, the conclusion holds. $\Box$
+Proposition 5.
 For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $s$ in $[m,4m+3]$.
@@
 Thus, Claim 1 holds for $i=1$.
-Now assume that $t$ is an integer with $t\geq 2$ such that $f_{1}(i),f_{2}(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $i$ with $11\leq i\leq t-$ 1. Then, by definition,
+Now assume that $t$ is an integer with $t\geq 2$ such that $f_{1}(i),f_{2}(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $i$ with $1\leq i\leq t-1$. Then, by definition,
 $\frac{2mi+i+m}{i\sqrt{4m+3}}=\sqrt{4m+3}-\frac{i+1}{\frac{(i+1)\sqrt{4m+3}}{2i+1}}-\frac{m-i}{\frac{i\sqrt{4m+3}}{2i-1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (37)
@@
 Thus, Claim 1 holds.
-Claim 2: For any integer $i$ with $-3\leq i\leq m$, $\frac{2m-i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
+Claim 2: For any integer $i$ with $3\leq i\leq m$, $\frac{2m-i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
 Let $g(i)=\frac{2m-i}{\sqrt{4m+3}}$. By Claim 1, $g(-3)=\frac{2m+3}{\sqrt{4m+3}}=f_{2}(m+1)$. By definition,
@@
 Hence Claim 2 follows.
-Claim 3: For any $s\in{\mathbb{N}}^{+}$ with $32m\leq s\leq 4m+$ 3, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
-By Lemma 4.1, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $33m+2\leq s\leq 4m+$ 3.
+Claim 3: For any $s\in{\mathbb{N}}^{+}$ with $2m\leq s\leq 4m+3$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
+By Lemma 4, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $3m+2\leq s\leq 4m+3$.
 Let $h(i)=\frac{2m+i}{\sqrt{4m+3}}$.
@@
 $h(i+2)=\frac{2m+2+i}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{(m+1-i)}{\frac{(m-i+1)\sqrt{4m+3}}{2\cdot(m-i)+1}}-\frac{i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (48)
-Thus, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $12m\leq s\leq 3m+$ 1. Hence Claim 3 holds. The result then follows from Claims 2 and 3. $\Box$
-We are now going to prove Theorem 1.5.
-Proof of Theorem 1.5: The necessity follows from (1). Now we prove the sufficiency.
+Thus, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $2m\leq s\leq 3m+1$. Hence Claim 3 holds. The result then follows from Claims 2 and 3. $\Box$
+We are now going to prove Theorem 3.
+Proof of Theorem 3: The necessity follows from Theorem 1. Now we prove the sufficiency.
 Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$.
-By Lemma 4.1 and Propositions 4.24.34.4 and 4.5, for each integer $s$ with $r_{0}\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})\subseteq{\mathscr{W}}_{s}(\sqrt{k})$ holds.
+By Lemma 4 and Propositions 234 and 5, for each integer $s$ with $r_{0}\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})\subseteq{\mathscr{W}}_{s}(\sqrt{k})$ holds.
 Note that
 $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}.$ (49)
-By Corollary 3.3, for any integer $s$ in $[r_{0},k]$, there exists a tree $T$ with $\Delta(T)=s$ and $\lambda(T)=\sqrt{k}$. $\Box$
+By Corollary 3, for any integer $s$ in $r_{0},k]$, there exists a tree $T$ with $\Delta(T)=s$ and $\rho(T)=\sqrt{k}$. $\Box$
 5 Further study
-Recall that $\Lambda$ is the set of $\lambda(T)$ over all trees $T$ of order at least $2$.
-Conjecture 1.4 asks an interesting question on the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$ for any given $\alpha\in\Lambda\setminus\{1\}$ and integer $r$ satisfying $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
-Another interesting problem is on the numbers contained in $\Lambda$.
+Recall that $\mathbb{P}$ is the set of $\rho(T)$ over all trees $T$ of order at least $2$.
+Problem 1 asks an interesting question on the existence of a tree $T$ with $\rho(T)=\alpha$ and $\Delta(T)=r$ for any given $\alpha\in\mathbb{P}\setminus\{1\}$ and integer $r$ satisfying $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
+Another interesting problem is on the numbers contained in $\mathbb{P}$.
 Let $\cal P$} be the set of real-rooted monic polynomials with integer coefficients.
@@
 a root of some polynomial in $\cal P$}) is an eigenvalue of some finite tree.
-Now let $\Lambda^{\prime}$ be the set of real numbers $\alpha$ such that $\alpha$ is the largest real root of some $P\in{\cal P}$.
-Problem 5.1.
-Is there a positive real number $c$ such that $\{\alpha\in\Lambda^{\prime}:\alpha\geq c\}\subseteq\Lambda$?
-It can be proved by applying Theorem 1.1 that, for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if $4|(p-q)$, $\sqrt{p+\sqrt{q}}\in\Lambda$ if $p\geq q$, and $\frac{\sqrt{(2p+1)+\sqrt{4q+1}}}{\sqrt{2}}\in\Lambda$ if $q\leq p^{2}+p$.
-As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ whenever $4|(p-q)$.
-Proposition 5.2.
-For any $p,q\in{\mathbb{N}}^{+}$ with $p\geq q$, if $p-q$ is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\left\{q,\frac{p-q}{4}\right\}$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
-Proof. Assume that $p-q=4d$. Let $r=\max\{q,d\}+1$ and $\alpha=\frac{\sqrt{p}+\sqrt{q}}{2}$. By definition,
+Now let $\mathbb{P}^{\prime}$ be the set of real numbers $\alpha$ such that $\alpha$ is the largest real root of some $P\in{\cal P}$.
+Problem 2.
+Is there a positive real number $c$ such that $\{\alpha\in\mathbb{P}^{\prime}:\alpha\geq c\}\subseteq\mathbb{P}$?
+It can be proved by applying Theorem 2 that, for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ if 4|(pq)4|(p-q), $\sqrt{p+\sqrt{q}}\in\mathbb{P}$ if $p\geq q$, and $\frac{\sqrt{(2p+1)+\sqrt{4q+1}}}{\sqrt{2}}\in\mathbb{P}$ if $q\leq p^{2}+p$.
+As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ whenever 4|(pq)4|(p-q).
+Proposition 6.
+For any $p,q\in{\mathbb{N}}^{+}$ with $p\geq q$, if pqp-q is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\left\{q,\frac{p-q}{4}\right\}$ and $\rho(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
+Proof. Assume that pq=4dp-q=4d. Let $r=\max\{q,d\}+1$ and $\alpha=\frac{\sqrt{p}+\sqrt{q}}{2}$. By definition,
 $\sqrt{q}=\frac{\sqrt{p}+\sqrt{q}}{2}-\frac{d}{\frac{\sqrt{p}+\sqrt{q}}{2}}\in{\mathscr{W}}_{r}(\alpha)$ (50)
@@
 $d\cdot\left(\frac{\sqrt{p}-\sqrt{q}}{2}\right)^{-1}=\frac{\frac{p-q}{4}}{\frac{\sqrt{p}-\sqrt{q}}{2}}=\frac{\sqrt{p}+\sqrt{q}}{2}=\alpha.$ (52)
-By Corollary 3.4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
-We wonder if $4|(p-q)$ is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\Lambda$.
-Problem 5.3.
-Is it true that for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if and only if $4\big{|}(p-q)$?
-Let $T$ be any tree with $\alpha=\lambda(T)$ and $r=\Delta(T)$.
-Then, (1) yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
+By Corollary 4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\rho(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
+We wonder if 4|(pq)4|(p-q) is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\mathbb{P}$.
+Problem 3.
+Is it true that for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ if and only if $4\big{|}(p-q)$?
+Let $T$ be any tree with $\alpha=\rho(T)$ and $r=\Delta(T)$.
+Then, Theorem 1 yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
 By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$.
-From the proving of Lemma 2.3, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
+From the proving of Lemma 1, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
 We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
-Problem 5.4.
-Let $\lambda\in\Lambda$ and $r\in{\mathbb{N}}^{+}$ with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?
+Problem 4.
+Let $\lambda\in\mathbb{P}$ and $r\in{\mathbb{N}}^{+}$ with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?
 Note that some similar work on Laplacian spectral radii of trees has been completed in [8].